### v1

# PEVLM: Parallel Encoding for Vision-Language Models

Abstract
Vision-Language Models (VLMs) have demonstrated strong performance in video-language tasks, yet their application to long video understanding remains constrained by the quadratic complexity of standard attention mechanisms. In this paper, we propose PEVLM, a parallel encoding strategy specifically designed to improve the prefill efficiency of VLMs without requiring model finetuning. PEVLM partitions the input into block-wise segments with a shared sink, preserves Full-Attention positional embeddings, and aligns attention weights to mimic Full-Attention distributions. This design reduces attention computation from $O((T\times N)^{2})$ to $O(T\times N)$ while maintaining high accuracy. Extensive experiments on the LongVideoBench benchmark show that PEVLM achieves up to 8.37% accuracy improvement over existing inference-efficient methods and delivers up to 7.47x speedup in attention computation and 40% reduction in end-to-end latency. Under strict latency constraints, PEVLM significantly outperforms baselines, raising accuracy from 23.26% to 61.03%. These results highlight PEVLM’s effectiveness for low-latency, long-context video understanding, making it well-suited for real-world applications such as autonomous driving.
1 Background
1.1 Challenges in Long Context Processing
In the realm of computer vision and natural language processing integration, Vision-Language Models (VLMs) have demonstrated remarkable performance in various multimodal tasks, particularly in video-text understanding and generation (alayrac2022flamingo, ; radford2021learning, ; vlm_performance_example, ). However, their application to long video processing remains computationally challenging due to the quadratic complexity of self-attention mechanisms (long_video_processing_challenges_example, ; vaswani2017attention, ). Specifically, VLMs usually extract frame-wise features and integrate textual context information, leading to a drastic increase in context length as each frame is tokenized into multiple embeddings.
The high computational complexity of VLMs presents a significant barrier to their practical use and deployment in real-world scenarios (li2022video, ). Take autonomous driving as an example: training advanced autonomous driving models requires extensive long video data, which in turn demands efficient data pre-processing and annotation techniques. Traditional manual annotation methods and classical annotation models are inefficient in handling the huge volume of long video data. Although VLMs offer the possibility of automated annotation, the computational cost of processing long videos is prohibitively high. Moreover, autonomous driving systems impose strict requirements on low latency execution and high accuracy due to the critical need for real-time decision making to ensure safety (zhou2022efficient, ). However, on-board computing resources are severely limited, and the latency introduced by long-context attention mechanisms in VLMs significantly degrades the response time of autonomous systems. This situation highlights the urgent challenge of improving the inference performance of the VLMs with limited latency while maintaining its accuracy.
In autonomous driving scenarios, the system typically does not generate lengthy CoT (chain-of-thought) analyses, but instead needs to output only short environmental features and final decisions. This results in the prefill stage accounting for a substantial proportion of the overall inference time, becoming the primary performance bottleneck. However, most existing methods for optimizing the prefill stage involve a trade-off between performance improvement and accuracy degradation. Given that autonomous driving and other safety-critical applications are highly intolerant of accuracy loss, these methods are not applicable. Therefore, our research focuses on search for novel strategies to optimize the prefill stage latency of VLMs with minimal impact on accuracy.
1.2 Existing Solutions and Limitations
A widely adopted technique in LLMs is the parallel-encoding mechanism (li2024focusllmpreciseunderstandinglong, ; liu2024star, ; ma2025blockattentionefficientprefilling, ; ratner2023parallelcontextwindowslarge, ; yen2024longcontextlanguagemodelingparallel, ; yang2025ape, ). In this approach, the input context is divided into blocks, and each block is encoded into Key-Value (KV) states independently. This means that tokens from different blocks do not attend to each other during encodin. During decoding, however, the model computes attention using the complete KVCache that contains all historical blocks.
In addition, several block-based methods have been proposed that introduce training procedures to further enhance model accuracy and performance (ma2025blockattentionefficientprefilling, ; yuan2025nativesparseattentionhardwarealigned, ). However, not all application scenarios have sufficient resources to support large-scale training. On the other hand, many platforms offering inference services cannot control users’ preferences for model usage, nor can they mandate further fine-tuning on existing models. This is especially true for autonomous driving models that have undergone continuous training and iteration over months or even years. Therefore, we prioritize finetune-free approaches for optimization.
The parallel-encoding mechanism omits the attention calculation between blocks, reduces the computational complexity from $O(L^{2})$ to $O(L)$. It also mitigates the problem of "lost in the middle" (Liu2023LostIT, ; yang2025ape, ). Several architectural improvements have been proposed to enhance the parallel-encoding mechanism:
  • Adaptive Parallel Encoding (APE) improves prefill efficiency by introducing shared prefixes, an attention temperature parameter, and a scaling factor to better align the distribution of parallel-encoding with that of sequential encoding (yang2025ape, ). Although these adjustments lead to improved accuracy, they are highly sensitive to length, quantity, and content, making them difficult to tune and deploy in practice.
  • Star Attention, introduces a two-phase block-sparse approximation that improves computational efficiency by sharding attention across multiple hosts while minimizing communication overhead and reducing computational complexity (liu2024star, ). Although primarily designed for distributed settings, its core algorithms are mathematically similar to APE when temperature and scaling are not applied.
  • Block-Attention applies position re-encoding to allow reuse of prefilled contexts, and maintains model accuracy through fine-tuning (guu2020retrieval, ). Without fine-tuning, its accuracy suffers due to the absence of a Sink Block, though the idea of retaining Full-Attention position encodings proves insightful.
Thus, a natural research question thus arises: Can such methods be directly applied to Vision-Language Models (VLMs) to accelerate inference and deployment pipelines?
2 Our Observations and Contributions
In Table 1, we evaluate several existing methods and find that parallel-encoding-based approaches are also effective in VLMs. These methods achieve competitive accuracy on LongVideoBench (longvideobench2024, ) without requiring any modifications. This observation is consistent with findings from APE in LLM scenarios (yang2025ape, ).
2.1 Observations Specific to VLM
Despite the similarities between LLMs and VLMs, there are important distinctions. As shown in Figure 1: (i) In VLMs, attention sinks are present not only in the initial system prompt, but also in the early video frames, which aligns with the findings of Kang et al. (kang2025toldvisualattentionsink, ). To achieve optimal performance, it may be necessary to include all frames with high attention scores in the sink. This requirement may vary slightly across models; for the model shown in the figure (Qwen2.5-VL), the optimal range appears to be around 8–12 frames. (ii) Existing parallel-encoding-based are primarily designed for RAG (Retrieval-Augmented Generation) tasks. In such tasks, reusing the KVCache from historical passages is crucial for efficiency, and the order of passages often does not matter. Consequently, position embeddings are typically reused across blocks.
However, video-understanding tasks differ significantly: Maintaining the correct temporal order between video frames is essential. Reusing position embeddings indiscriminately across blocks can impair temporal comprehension of the model.
(a) Full-Attention
(b) Reuse Position
(c) Keep Full-Attention Position Embedding
Figure 1: Comparison of Attention Weight Distributions across Video Contexts. The top row shows the attention weights for all tokens, while the bottom row shows those for the video tokens only. (a) Full-Attention: The textual tokens and the initial video frames receive the highest attention weights, whereas the remaining video frames are assigned relatively lower attention weights. (b) Parallel Encoding with Reused Position Embeddings: The overall attention weights distribution resembles Full-Attention, with higher attention weights for the textual tokens and initial video frames. While compared to Full-Attention, the initial video frames receive even higher attention weights, and the remaining video frames exhibit more uniform and lower attention weights. (c) Parallel Encoding with Full-Attention Position Embedding: The attention pattern remains similar to (a) and (b), highlighting the textual tokens and the initial video frames. However, the attention weight distribution within the video frames more closely aligns with that of Full-Attention.
Therefore, as illustrated in Figure 1, we perform an attention weight analysis for VLMs and make two key findings: (i) Initial segments of both text and video receive high attention weights, emphasizing the need to include early video frames when selecting sink tokens. (ii) Using Full-Attention position embeddings (rather than reusing position embeddings across blocks) yields attention distributions more similar to Full-Attention, leading to better alignment with the model’s intended behavior. These insights guide the design of VLM-specific adaptations of parallel-encoding that preserve both efficiency and contextual integrity.
2.2 Our Contributions
Basing on these observations, we have conducted a series of studies and experiments. And proposed a novel attention mechanism for VLMs named Parallel Encoding for Vision-Language Models (PEVLM), which is a finetune-free method specifically designed to speed up prefill of VLM-based long-video processing. By alleviating computational bottlenecks in long video prefill, PEVLM facilitates the efficient deployment of VLMs in various practical application scenarios on both clouds and terminal equipments.
Our contributions involve:
  • We introduce PEVLM, a blockwise parallelization strategy for VLMs. PEVLM reduces the computational complexity of the attention mechanism from $O((T\times N)^{2})$ to $O(T\times N)$, which leading to a significant improvement in prefill efficiency.
  • We recover the accuracy of PEVLM with three alignment steps: (i) Segmenting contexts into blocks by video frames rather than tokens. (ii) Using the system promots and the initial frames of the video as Sink Block for all blocks to avoid the duplication of abnormal distribution of initial tokens. (iii) Preserving the position embeddings of Full-Attention instead of resuing position embeddings across blocks. This suggests that PEVLM serves as a performance optimization mechanism rather than endowing the model with the capability to process context lengths that exceed its inherent capacity.
  • We empirically show that (i) PEVLM gets better accuracy than Block-Attention, APE and Star Attention in long video processing tasks, achieving accuracy improvements ranging from 0.68% to 8.37% across different models. (ii) For 100k token contexts, the attention layer achieves a 7.47× speedup, while the end-to-end processing time is reduced by 40% without compromising generation quality. This highlights the practical viability and superiority of PEVLM in cloud deployment scenarios. (iii) Under fixed latency constraints, PEVLM increases accuracy from 23.26% to 61.03%, showcasing its critical value for latency sensitive applications (e.g., autonomous driving).
3 PEVLM
PEVLM is an advanced attention mechanism designed to improve computational efficiency of prefill while maintaining global contextual consistency. This approach draws from multiple existing techniques to address computational bottlenecks commonly encountered in long video processing.
3.1 Standard Attention Mechanism
In a standard Softmax attention, we attend the query to all past KV states using the following formula:
$O=Softmax(\frac{QK^{T}}{\sqrt{d}})V\ \ \ Q\in\mathbb{R}^{n\times d}\ \ \ K,V\in\mathbb{R}^{m\times d}$ (1)
where $Q$ is the query state, and $K$ and $V$ denote the key and value states, respectively.
Attention Sink. First discovered by StreamingLLM (mit2023streamingllm, ), that one or more tokens at the start position of an LLM receive higher attention scores. Some VLM-related studies also show that the attention sink phenomenon exists in VLMs but is more distributed among image tokens (kang2025toldvisualattentionsink, ).
Position Embedding. To effectively process sequential input, LLMs require position embeddings, such as absolute position embeddings (devlin-etal-2019-bert, )(vaswani2017attention, ) and relative position embeddings (press2022trainshorttestlong, )(SU2024127063, ). For VLMs, contextual positional information is more critical. In particular, for some VLM models like Qwen2.5-VL, their position embeddings also incorporate spatio-temporal information between video frames, making position embeddings increasingly essential for VLMs.
Computational Complexity. In standard Softmax attention mechanisms, the total OPs is
$OPs_{Attn}=2HL^{2},$ (2)
where $L$ denotes the context length and $H$ is the hidden size. So, the computational complexity is $O(L^{2})$. As $L$ increases, the computational cost increases significantly, restricting the inference speed and scalability. It is even worse for VLMs. Given a video with $T$ frames, where each frame is represented by $N$ tokens (typically, $N$ ranges between 200 and 600), the resulting context length $L$ scales to $(T\times N)$ and complexity to $O((T\times N)^{2})$, making standard attention mechanisms impractical for long video understanding tasks.
3.2 Parallel Encoding
3.2.1 Partitioning Strategy
Figure 2: The PEVLM Masks.
APE (yang2025ape, ) and Star-Attention (liu2024star, ) have demonstrated that sharing a common attention sink across blocks significantly enhances model accuracy. Building on recent advances in efficient attention mechanisms and long-context modeling (mit2023streamingllm, ; liu2024star, ; xiao2023efficient, ; yang2025ape, ), we propose a Sink Block mechanism to improve inference quality. As illustrated in Figure 2, our partitioning strategy consists of three core components:
  • Sink Block: The initial text tokens (e.g., BOS token and system prompts) and the first several video frames are grouped into a dedicated Sink Block.
  • Context Blocks: The remaining video content is uniformly partitioned into Context Blocks by frames to reduce computational overhead and enable parallelization.
  • Question Block: The text tokens that follow the video input are left unsegmented as Quesion Block.
Since the information within each video frame represents spatial content at a specific moment, while information across frames reflects temporal dynamics, videos naturally possess structural boundaries. If the video are divided into sink and context blocks by token count, this may compromise the spatial integrity of the boundary frames shared between adjacent blocks. So we divide the video by frame.
In addition, as analyzed in Section 2.1, the number of frames selected for the Sink Block depends primarily on the distribution of attention weights. We conduct further empirical studies and analyses on this in Section 5.2. On the other hand, the number of frames per Context Block involves a trade-off between accuracy and computational efficiency, which we investigate and analyze in Section 5.3.
3.2.2 Formulations
For different blocks, PEVLM applies distinct computation methods to ensure efficient processing and information integration:
$\text{Attn}_{s}\ =f(Q_{s},K_{s},V_{s}),$ (3)
$\text{Attn}_{c_{i}}=f(Q_{c_{i}},K_{s+c_{0}+\dots+c_{i-1}},V_{s+c_{0}+\dots+c_{i-1}}),$ (4)
$\text{Attn}_{q}\ =f(Q_{q},K_{s+c_{\text{all}}+q},V_{s+c_{\text{all}}+q}),$ (5)
where $s$ and $q$ are the Sink Block and question block, $c_{i}$ is the ith Context Blocks; $Attn_{s}$, $Attn_{c_{i}}$ and $Attn_{q}$ denote the attention of the $s$, $c_{i}$ and $q$. $f(\cdot)$ represents the standard Softmax attention in Equation (1), and $Q$, $K$ and $V$ denote the query, key, and value matrices.
The total OPs of PEVLM is
$OPs_{PEVLMAttn}=OPs_{SinkAttn}+N\times OPs_{BlockAttn}+OPs_{QuestAttn},$ (6)
$OPs_{SinkAttn}=2HS^{2},$ (7)
$OPs_{BlocksAttn}=2HB(S+B),$ (8)
$OPs_{QuestAttn}=2HQL,$ (9)
where $S$ denotes the sink size, $N$ denotes the block number, $B$ denotes the block size, $Q$ denotes the query size and $H$ is the hidden size. $L$ is the total token number and $L=S+N\times B+Q$.
After simplification:
$OPs_{PEVLMAttn}=2H(S^{2}+Q^{2}+NB^{2}+QS+NQB+NSB).$ (10)
As B is a static number, the computational complexity of PEVLM is $O(L)$.
3.2.3 Position Encoding in VLMs
Figure 3: Position Encoding in PEVLM
As described in Section 2.1, similar to the case in LLMs, applying parallel-encoding-based methods in VLMs suffers from the misalignment of attention weight distributions when compared to Full-Attention. However, introducing additional hyperparameters to adjust attention weights for better alignment is not ideal for efficient deployment. Therefore, inspired by Block Attention (ma2025blockattentionefficientprefilling, ), we retain Full-Attention Position Embedding as illustrated in Figure 3.
Due to maintaining the original position embedding, PEVLM cannot enable the model to handle longer video contexts beyond its inherent capabilities like some other parallel-encoding-based methods. For example, LongVILA-7B-256f can originally process 256 frames at most. Although using PEVLM can accelerate the inference speed, it still processes 256 frames at most. This is not a problem as our goal is to accelerate the prefill.
4 Experiments
In this section, we test PEVLM from multiple perspectives against several different VLMs. The primary objective is to evaluate the accuracy, computational efficiency, and real-world applicability of PEVLM for long-video processing.
4.1 Accuracy Evaluation
The main objective of this experiment is to analyze the impact of PEVLM on accuracy in long-video understanding tasks.
4.1.1 Setup
We have chosen the LongVideoBench (longvideobench2024, ) as a benchmark, which is designed to comprehensively evaluate large multimodal models on long-term multimodal understanding. We have carried out experiments using LLaVA-Video (zhang2024videoinstructiontuningsynthetic, ), Qwen2.5-VL (qwen2.5-VL, ), and LongVILA (chen2024longvilascalinglongcontextvisual, ) as test models. To minimize any adverse effects on model accuracy, all evaluations were performed using bf16 precision.
As shown in Table 1, we compared the accuracy of Full-Attention, Block-Attention, APE, Star-Attention, and PEVLM in the LongVideoBench dataset. Since the other methods are finetune-free, we did not finetune the weights in the Block Attention test. For APE and Star-Attention, we tested two scenarios: (i) reusing position encoding across all blocks; (ii) retaining the position encoding of Full-Attention. For PEVLM, we set the sink size and block size to about 16 frames, which corresponds to approximately 4k tokens for Qwen2.5-VL and LongVILA, and about 3k tokens for LLaVA-Video.
4.1.2 Results
Table 1: LongVideoBench Accuracy
model method sink(anchor)_size block_size Correct $\uparrow$
Full Attn - - 60.43%
Block Attn (wo finetune) 0 4096 54.23%
APE 256 4096 NaN
Qwen2.5-VL- APE (retain pos) 256 4096 60.21%
7B-Instruct Star Attn 4096 4096 NaN
Star Attn (retain pos) 4096 4096 61.41%
PEVLM sys+16frames 16frames 62.23%
Full Attn - - 52.21%
Block Attn (wo finetune) 0 4096 51.76%
APE 256 4096 51.68%
LongVILA- APE (retain pos) 256 4096 53.18%
7B-256f Star Attn 4096 4096 46.75%
Star Attn (retain pos) 4096 4096 53.63%
PEVLM sys+16frames 16frames 53.70%
Full Attn - - 59.99%
Block Attn (wo finetune) 0 4096 55.50%
APE 256 4096 58.41%
LLaVA-Video- APE (retain pos) 256 4096 58.49%
7B-Qwen2 (128) Star Attn 4096 4096 58.71%
Star Attn (retain pos) 4096 4096 58.49%
PEVLM sys+16frames 16frames 59.01%
The experiments demonstrate that PEVLM achieves an accuracy comparable to or even exceeding that of Full-Attention across the three tested models. Both APE and Star-Attention can also approach or surpass Full-Attention’s accuracy when retaining Full-Attention’s position encoding. However, for Block-Attention, due to the lack of resources for alignment fine-tuning, its accuracy shows a significant decline compared to Full-Attention.
Without keeping the position encoding of Full-Attention, the output of the Qwen2.5-VL-7B-Instruct model exhibited severe output garbling and semantic confusion, and the LongVILA model also experienced noticeable performance degradation. However, the LLaVA-Video model showed insensitivity to changes in position encoding, with negligible accuracy variation observed regardless of whether Full-Attention position encoding was retained.
4.2 Performance Evaluation
In this section, we evaluate the performance of PEVLM. To facilitate experimentation, we built PEVLM on the SGLang framework, which is a fast serving framework for large language models and vision-language models, widely used in cloud production deployment. And to better reflect practical deployment scenarios, we designed the experiments from two perspectives. (i) We first test with workloads of approximately 100k tokens to assess the acceleration performance of PEVLM in cloud computing environments. (ii) Furthermore, we evaluate PEVLM on the LongVideoBench dataset under a given latency constraint, in order to measure its effectiveness in edge scenarios where both computational resources and latency are limited.
4.2.1 Setup
Since the LLM component in VLM models remains largely consistent across different implementations, performance differences arise primarily from hyperparameter configurations rather than model weights. Consequently, we conducted computational efficiency evaluations on the Qwen2.5-VL model. The primary objective of this experiment is to assess the improvements in computational efficiency introduced by PEVLM in computational efficiency. All experiments were carried out on an H20-96G machine. Given that PEVLM primarily optimizes the attention mechanism within the LLM, we measured the execution time of both the entire LLM module and its individual attention layers separately.
Figure 4: We evaluated the performance of PEVLM under different configurations of (sink_size / block_size). Using the Qwen2.5-VL-7B-Instruct model, we tested its LLM and Attention time on a video dataset with an average duration of 10 minutes (approximately 106k tokens per sample). "Attn Time" designates the Attention Computation time, whereas "Other Time" signifies the aggregate of all remaining computational costs within the LLM excluding Attention Computation. The "Attn SpeedUP" represents the performance acceleration of PEVLM relative to the "Full Attn". The "LLM SpeedUP" denotes the overall performance acceleration of the LLM in PEVLM compared to the "Full Attn".
4.2.2 Results
As illustrated in Figure 4, the end-to-end inference speed of the LLM is improved by 2.58× to 3.14×. Notably, the attention module achieves a speedup ranging from 7.79× to 33.81×, while the execution time of non-attention components remains largely unaffected. As the block size or sink size increases, the acceleration effect of the attention module decreases significantly. This observation is consistent with the analysis presented in the Appendix. Moreover, because non-attention components constitute a substantial portion of the total runtime in PEVLM, the overall end-to-end speedup of the LLM moderately decreases from 3.14× to 2.58×.
4.3 Accuracy with Limited Latency
One key challenge in deploying VLMs on edge devices is the limited computational resources and strict latency requirements. To evaluate the acceleration performance of PEVLM under this scenario, we designed a simulation accordingly. A simplified experimental setup was implemented to assess the efficiency of PEVLM in edge deployment settings.
4.3.1 Setup
We use LongVideoBench as the evaluation benchmark and introduce a latency constraint during inference: any sample whose inference time exceeds the predefined latency threshold is considered a failure. The experiments are conducted on the H20-96B platform. We select Qwen2.5-VL-7B-Instruct with Full-Attention as the baseline, and compare several methods, including Full-Attention, Block Attention, APE, Star Attention, PEVLM, and using smaller models (Qwen2.5-VL-3B-Instruct). Since the other methods are finetune-free, we did not finetune the weights in the Block Attention test. Although position encoding does not affect inference lantency, reusing position encoding performs extremely poorly accuracy on the Qwen2.5-VL model. Therefore, in the APE and Star Attention tests, we adopted the method of retaining Full-Attention’s position encoding.
4.3.2 Results
Table 2: LongVideoBench Accuracy with Limited Latency
method sink(anchor) block No Limit $\uparrow$ 40s $\uparrow$ 30s $\uparrow$ 20s $\uparrow$
Full Attn - - 60.43% 59.69% 27.08% 23.26%
Full Attn (3B) - - 55.05% 55.05% 55.05% 24.38%
Block Attn 0 4096 54.00% 54.00% 53.78% 53.55%
APE 256 4096 59.91% 59.91% 59.91% 59.61%
Star Attn 4096 4096 61.41% 61.41% 61.41% 28.20%
PEVLM sys+2frames 16frames 61.18% 61.18% 61.18% 61.03%
PEVLM sys+16frames 16frames 62.23% 62.23% 62.23% 60.28%
As shown in Table 2, the accuracy of Full-Attention drops significantly when the latency constraint is reduced from 40s to 30s, while the accuracy of Full-Attention 3B also declines sharply when the latency constraint is tightened from 30s to 20s. Due to its larger sink and block sizes, StarAttn gets more OPs and exhibits a noticeable accuracy drop between 30s and 20s. In contrast, Block Attention, APE, and PEVLM maintain relatively high accuracy even at 20s. Notably, although larger sink and block sizes (e.g., sys+16frames, block=16frames) achieve higher accuracy under looser latency constraints, the smaller sink size (sink=sys+2frames, block=16frames) delivers better accuracy at the 20s latency limit due to its lower latency.
5 How does each component in PEVLM contribute to the performance?
5.1 Split The Video by Frames instead of Tokens
Due to the inherent nature of videos being composed of multiple frames, where intra-frame information captures spatial details at a single time point, and inter-frame information represents temporal dynamics, ignoring this structure and arbitrarily dividing blocks based on token count may compromise the spatial integrity at critical temporal boundaries. Therefore, we choose to divide the video into blocks based on frames. This design is supported by the comparison between Star Attention with Full-Attention positional encoding and PEVLM, as shown in Table 1. Across different models, frame-based block partitioning consistently provides accuracy gains, albeit varying in magnitude, over token-based partitioning.
5.2 Sink Size
As analyzed in Section 3.2.2, the inference latency of PEVLM increases quadratically with the sink size, which is consistent with the experimental results shown in Figure 4.
Furthermore, we investigate the impact of sink size on model accuracy. As observed in Section 2.1, it is essential to involve the initial frames of the video in the sink. This observation is validated in Figure 5, where incorporating the initial frames into the sink leads to varying degrees of accuracy improvement across different models.
Moreover, as shown in Figure 5, the accuracy of Qwen2.5-VL and LLaVA-Video shows an upward trend as the sink size increases, though the improvement remains marginal. In contrast, LongVILA exhibits fluctuations in accuracy with varying sink sizes but does not show a consistent upward or downward trend. However, because of the quadratic increase in latency, using longer sink sizes is not favorable for performance. Therefore, it may be necessary to conduct a simple test prior to deployment to determine an appropriate sink size. For the three models shown in Figure 5, setting the sink size between 14 and 18 already yields strong accuracy. If latency constraints are also taken into account, this could lead to a trade-off between performance and accuracy. For example, as demonstrated in the results of Section 4.3, under a 20-second latency constraint, using only sys+2frames as the Sink Block achieves better accuracy than using sys+16frames.
Figure 5: The impact of sink/block sizes on accuracy
5.3 Block Size
As analyzed in Section 3.2.2, the inference latency of PEVLM also increases quadratically with the block size as well as the sink size, which is consistent with the experimental results shown in Figure 4. However, the block size has a more significant impact on performance because the coefficient of block-size’s quadratic term in Equation 10 is much larger than that of the sink size.
We investigate the impact of block size on model accuracy here. As shown in Figure 5, the model accuracy exhibits an upward trend as the block size increases, similar to sink size. However, due to the quadratic growth in latency, a trade-off arises between computational efficiency and accuracy.
5.4 Position Encoding
As analyzed in Section 2.1, using Full-Attention position embeddings, rather than reusing position embeddings across blocks, should yield attention distributions that more closely resemble those of Full-Attention, thereby aligning better with the model’s intended behavior. This is validated by the experiments presented in Section 4.1. As shown in Table 1, for several parallel-encoding-based methods (Block-Attention, APE, Star Attention, and PEVLM), using Full-Attention position embeddings significantly improves accuracy in both Qwen2.5-VL and LongVILA. In contrast, for the LLaVA-Video model, the benefit is marginal, likely because its attention weight distribution remains relatively stable even when using parallel encoding.
6 Conclusion
This study investigates the potential of parallel-encoding-based methods in VLMs, which can reduce the computational complexity of attention mechanisms for fast inference and reuse positions for long-context processing but often result in performance degradation. To address this issue, we propose PEVLM, a training-free method designed to enable accurate, fast, and long-context video processing. PEVLM achieves this by aligning the attention weight distribution of parallel encoding with Full-Attention through two key steps: selecting an appropriate sink size and block size, and utilizing Full-Attention position embeddings. Empirically, we demonstrate that PEVLM enhances both accuracy and efficiency in long video processing tasks while successfully scaling to process hundreds of chunks in parallel across various settings.
7 Future Work
Although PEVLM has demonstrated substantial improvements in computational efficiency and reasoning performance for long video processing tasks, there remains ample opportunity for further research and enhancement. In future work, we plan to explore the following directions.
In autonomous driving scenarios, a common strategy is to use several historical frames as input to make decisions for the next action. Due to the computational nature of the attention mechanism, most current approaches require re-computation of all historical frames as the temporal step advances. Building upon the current PEVLM framework, we propose integrating the rotating positional embedding method from Block-Attention (ma2025blockattentionefficientprefilling, ), which enables the reuse of the KVCache from all historical frames preceding the current timestep. This approach has the potential to yield an order-of-magnitude improvement in efficiency for such applications.
Many of the existing methods based on parallel-encoding, including PEVLM, can achieve accuracy that exceeds Full-Attention. This phenomenon is attributed to the fact that parallel-encoding-based methods can mitigate the "lost in the middle" problem, thereby enabling them to achieve accuracy surpassing Full-Attention when processing long contexts (yang2025ape, ). The reason why parallel-encoding can mitigate the "lost in the middle" issue is hypothesized to be due to its shortening of the distance between the context and the query. However, more concrete evidence is needed to support this claim.

### v2

@@
 Abstract
-Vision-Language Models (VLMs) have demonstrated strong performance in video-language tasks, yet their application to long video understanding remains constrained by the quadratic complexity of standard attention mechanisms.
-In this paper, we propose PEVLM, a parallel encoding strategy specifically designed to improve the prefill efficiency of VLMs without requiring model finetuning.
-PEVLM partitions the input into block-wise segments with a shared sink, preserves Full-Attention positional embeddings, and aligns attention weights to mimic Full-Attention distributions.
-This design reduces attention computation from $O((T\times N)^{2})$ to $O(T\times N)$ while maintaining high accuracy.
-Extensive experiments on the LongVideoBench benchmark show that PEVLM achieves up to 8.37% accuracy improvement over existing inference-efficient methods and delivers up to 7.47x speedup in attention computation and 40% reduction in end-to-end latency.
-Under strict latency constraints, PEVLM significantly outperforms baselines, raising accuracy from 23.26% to 61.03%.
-These results highlight PEVLM’s effectiveness for low-latency, long-context video understanding, making it well-suited for real-world applications such as autonomous driving.
-1 Background
-1.1 Challenges in Long Context Processing
-In the realm of computer vision and natural language processing integration, Vision-Language Models (VLMs) have demonstrated remarkable performance in various multimodal tasks, particularly in video-text understanding and generation (alayrac2022flamingo, ; radford2021learning, ; vlm_performance_example, ).
-However, their application to long video processing remains computationally challenging due to the quadratic complexity of self-attention mechanisms (long_video_processing_challenges_example, ; vaswani2017attention, ).
-Specifically, VLMs usually extract frame-wise features and integrate textual context information, leading to a drastic increase in context length as each frame is tokenized into multiple embeddings.
-The high computational complexity of VLMs presents a significant barrier to their practical use and deployment in real-world scenarios (li2022video, ).
-Take autonomous driving as an example: training advanced autonomous driving models requires extensive long video data, which in turn demands efficient data pre-processing and annotation techniques.
-Traditional manual annotation methods and classical annotation models are inefficient in handling the huge volume of long video data.
-Although VLMs offer the possibility of automated annotation, the computational cost of processing long videos is prohibitively high.
-Moreover, autonomous driving systems impose strict requirements on low latency execution and high accuracy due to the critical need for real-time decision making to ensure safety (zhou2022efficient, ).
-However, on-board computing resources are severely limited, and the latency introduced by long-context attention mechanisms in VLMs significantly degrades the response time of autonomous systems.
-This situation highlights the urgent challenge of improving the inference performance of the VLMs with limited latency while maintaining its accuracy.
-In autonomous driving scenarios, the system typically does not generate lengthy CoT (chain-of-thought) analyses, but instead needs to output only short environmental features and final decisions.
-This results in the prefill stage accounting for a substantial proportion of the overall inference time, becoming the primary performance bottleneck.
-However, most existing methods for optimizing the prefill stage involve a trade-off between performance improvement and accuracy degradation.
-Given that autonomous driving and other safety-critical applications are highly intolerant of accuracy loss, these methods are not applicable.
-Therefore, our research focuses on search for novel strategies to optimize the prefill stage latency of VLMs with minimal impact on accuracy.
-1.2 Existing Solutions and Limitations
-A widely adopted technique in LLMs is the parallel-encoding mechanism (li2024focusllmpreciseunderstandinglong, ; liu2024star, ; ma2025blockattentionefficientprefilling, ; ratner2023parallelcontextwindowslarge, ; yen2024longcontextlanguagemodelingparallel, ; yang2025ape, ).
-In this approach, the input context is divided into blocks, and each block is encoded into Key-Value (KV) states independently.
-This means that tokens from different blocks do not attend to each other during encodin.
-During decoding, however, the model computes attention using the complete KVCache that contains all historical blocks.
-In addition, several block-based methods have been proposed that introduce training procedures to further enhance model accuracy and performance (ma2025blockattentionefficientprefilling, ; yuan2025nativesparseattentionhardwarealigned, ).
-However, not all application scenarios have sufficient resources to support large-scale training.
-On the other hand, many platforms offering inference services cannot control users’ preferences for model usage, nor can they mandate further fine-tuning on existing models.
-This is especially true for autonomous driving models that have undergone continuous training and iteration over months or even years.
-Therefore, we prioritize finetune-free approaches for optimization.
-The parallel-encoding mechanism omits the attention calculation between blocks, reduces the computational complexity from $O(L^{2})$ to $O(L)$.
-It also mitigates the problem of "lost in the middle" (Liu2023LostIT, ; yang2025ape, ).
-Several architectural improvements have been proposed to enhance the parallel-encoding mechanism:
-  • Adaptive Parallel Encoding (APE) improves prefill efficiency by introducing shared prefixes, an attention temperature parameter, and a scaling factor to better align the distribution of parallel-encoding with that of sequential encoding (yang2025ape, ).
-Although these adjustments lead to improved accuracy, they are highly sensitive to length, quantity, and content, making them difficult to tune and deploy in practice.
-  • Star Attention, introduces a two-phase block-sparse approximation that improves computational efficiency by sharding attention across multiple hosts while minimizing communication overhead and reducing computational complexity (liu2024star, ).
-Although primarily designed for distributed settings, its core algorithms are mathematically similar to APE when temperature and scaling are not applied.
-  • Block-Attention applies position re-encoding to allow reuse of prefilled contexts, and maintains model accuracy through fine-tuning (guu2020retrieval, ).
-Without fine-tuning, its accuracy suffers due to the absence of a Sink Block, though the idea of retaining Full-Attention position encodings proves insightful.
-Thus, a natural research question thus arises: Can such methods be directly applied to Vision-Language Models (VLMs) to accelerate inference and deployment pipelines?
-2 Our Observations and Contributions
-In Table 1, we evaluate several existing methods and find that parallel-encoding-based approaches are also effective in VLMs.
-These methods achieve competitive accuracy on LongVideoBench (longvideobench2024, ) without requiring any modifications.
-This observation is consistent with findings from APE in LLM scenarios (yang2025ape, ).
-2.1 Observations Specific to VLM
-Despite the similarities between LLMs and VLMs, there are important distinctions.
-As shown in Figure 1: (i) In VLMs, attention sinks are present not only in the initial system prompt, but also in the early video frames, which aligns with the findings of Kang et al.
-(kang2025toldvisualattentionsink, ).
-To achieve optimal performance, it may be necessary to include all frames with high attention scores in the sink.
-This requirement may vary slightly across models; for the model shown in the figure (Qwen2.5-VL), the optimal range appears to be around 8–12 frames.
-(ii) Existing parallel-encoding-based are primarily designed for RAG (Retrieval-Augmented Generation) tasks.
-In such tasks, reusing the KVCache from historical passages is crucial for efficiency, and the order of passages often does not matter.
-Consequently, position embeddings are typically reused across blocks.
-However, video-understanding tasks differ significantly: Maintaining the correct temporal order between video frames is essential.
-Reusing position embeddings indiscriminately across blocks can impair temporal comprehension of the model.
+Vision-Language Models (VLMs) have demonstrated strong capabilities in multimodal understanding and generation tasks.
+However, their application to long video understanding remains hindered by the quadratic complexity of standard attention mechanisms.
+In this work, we introduce PEVLM, a fine-tuning-free parallel encoding method designed to enhance the prefilling efficiency of VLMs in long video scenarios.
+PEVLM partitions the input video into context blocks with a shared sink block, while preserving sequential position embeddings to align the attention weight distribution with that of Full-Attention.
+This design reduces attention complexity from $O((T\times N)^{2})$ to $O(T\times N)$ where $T$ is the number of frames and $N$ the number of tokens per frame, without sacrificing accuracy.
+Extensive experiments across multiple state-of-the-art models and benchmarks demonstrate that PEVLM consistently outperforms existing parallel encoding approaches, achieving up to 7.47x speedup in attention computation and reducing end-to-end latency by 40%.
+Remarkably, PEVLM not only maintains high accuracy, but in some settings even surpasses Full-Attention performance.
+Under strict latency constraints, it achieves substantial gains, improving accuracy from 23.26% to 61.03%.
+These results underscore the effectiveness of PEVLM for low-latency, long-context video understanding, making it a promising solution for real-world applications.
+Introduction
+In recent years, Vision-Language Models (VLMs) have emerged as a central research focus at the intersection of computer vision and natural language processing, demonstrating impressive performance in multimodal understanding and generation tasks (Alayrac et al.
+2022; Li et al.
+2020; Radford et al.
+2021; Zhu et al.
+2023).
+As the capabilities of these models continue to strengthen, their application scenarios have also expanded, including robotics (Black, Lee et al.
+2024; Cheang, Huang et al.
+2024; Prasad, Singh et al.
+2024), autonomous driving (Gao, Lin et al.
+2024; Hu, Wu et al.
+2023; Wang, Zhao et al.
+2024), and healthcare (Liu, Zhang et al.
+2024).
+Correspondingly, the length of the videos has also increased significantly.
+However, applying VLMs to long-video data introduces a significant computational challenge: The quadratic complexity of transformer-based attention mechanisms causes computation to scale rapidly with the input sequence length (Beltagy, Peters, and Cohan 2020; Vaswani et al.
+2017), especially during the prefilling stage, where all input tokens must be processed before generating any output.
+This severely limits the scalability of VLMs in long-video understanding tasks.
+In practical scenarios where the entire video must be encoded before decoding begins—such as video captioning, retrieval, or reasoning—this inefficiency becomes a critical bottleneck.
+Therefore, optimizing the prefilling phase is essential to efficiently deploy VLMs on long video tasks.
+A widely adopted technique in LLMs to speed up long context prefilling is the parallel encoding mechanism (Li et al.
+2024b; Liu et al.
+2024; Ma, Wang, and Tian 2025; Ratner et al.
+2023; Yang, Chen, and Chen 2025; Yen, Gao, and Chen 2024; Lu et al.
+2024).
+In these approaches, the input context is divided into blocks, and each block is encoded into Key-Value (KV) states independently.
+This means that tokens from different blocks do not attend to each other.
+Since the parallel encoding mechanism omits attention computation across blocks, it reduces the computational complexity from $O(L^{2})$ to $O(L)$.
+The parallel encoding mechanism also mitigates the problem of "lost in the middle" (Liu et al.
+2023), and in long-context scenarios, it can even achieve accuracy surpassing that of Full-Attention (Yang, Chen, and Chen 2025).
+Several state-of-the-art architectural improvements have been proposed to enhance the parallel encoding mechanism.
+Adaptive Parallel Encoding (APE) improves prefilling efficiency by introducing shared prefixes, an attention temperature parameter, and a scaling factor to better align the distribution of attention weight in parallel encoding with that of sequential encoding.
+Although these adjustments improve accuracy, they are highly sensitive to length, quantity, and content, making them difficult to tune and deploy in practice (Yang, Chen, and Chen 2025).
+Star Attention extends ring attention by incorporating a parallel encoding strategy, which eliminates communication overhead during the prefilling stage and reduces overall computational complexity (Liu et al.
+2024).
+Another notable distinction of Star Attention is the use of a larger sink size, which is equal to the block size.
+Block-Attention & TurboRAG share the same idea: build on parallel encoding, apply position re-encoding and fint-tune to improve accuracy (Guu et al.
+2020; Lu et al.
+2024).
+However, without fine-tune, the accuracy of both suffers for the absence of a Sink Block.
+Nevertheless, the idea of preserving the sequential position embeddings for each block, rather than reusing a shared one, offers a valuable and insightful idea.
+In addition, there are several widely discussed methods such as Native Sparse Attention (Yuan et al.
+2025) from Deepseek and MoBA (Lu et al.
+2025), which adopt similar algorithmic principles and achieve acceleration effects comparable to parallel encoding.
+However, both approaches involve modifications to the model architecture and thus require retraining from scratch.
+This poses a challenge for many inference platforms, which typically do not have control over users’ model preferences.
+The issue is particularly pronounced in domains such as autonomous driving, where a reliable models are often continuously trained and iterated over years.
+Consequently, we focus on finetune-free optimization methods to ensure compatibility and ease of deployment.
+Given all of these, a key question naturally arises: Can existing methods be directly applied to VLMs to accelerate inference and deployment pipelines?
+In Table 1, we evaluate several models with existing parallel encoding methods on LongVideoBench (Zhang and Wang 2024), VideoMME (Fu et al.
+2024), EgoSchema (Mangalam, Akshulakov, and Malik 2023) and MVBench (Li et al.
+2024a) benchmarks.
+Although existing parallel encoding methods achieve some effectiveness in shorter videos, we observe varying degrees of accuracy degradation across different models on longer videos, particularly for Qwen2.5-VL.
+To investigate this issue, we analyze the attention weight distributions under parallel encoding and identify a misalignment between Full-Attention and parallel encoding.
+Existing methods developed for LLMs fail to effectively address this misalignment in VLMs.
+Building upon this observation.
+We conducted a series of studies and experiments and ultimately propose a novel attention mechanism for VLMs, termed Parallel Encoding for Vision-Language Models (PEVLM).
+PEVLM is a fine-tuning-free method specifically designed to accelerate the prefilling stage in long-video processing with VLMs.
+Our contributions involve:
+  • We systematically analyze the distribution characteristics of attention weights under parallel encoding in VLMs.
+Our observations reveal a key reason for the misalignment between the attention weight distributions of parallel encoding and Full-Attention: the reuse of position embeddings across blocks leads to the loss of critical video information.
+  • We propose PEVLM to recover the accuracy of parallel encoding by applying three targeted alignment steps: (i) Segmenting contexts into blocks by video frames rather than tokens.
+(ii) Using the system promots and the initial frames of the video as Sink Block for all blocks to avoid the duplication of abnormal distribution of initial tokens.
+(iii) Preserving the sequential position embeddings instead of resuing position embeddings across blocks.
+PEVLM achieves superior accuracy compared to other parallel-encoding-based methods on long video processing tasks, with up to a 9.73% improvement across different models.
+  • PEVLM reduces the computational complexity of the attention mechanism from $O((T\times N)^{2})$ to $O(T\times N)$, which leading to a significant improvement in prefilling efficiency.
+(i) For 100k token contexts prefilling, the attention layer achieves a 7.47× speedup, while the end-to-end 2.58x speedup without compromising generation quality.
+This highlights the practical viability and superiority of PEVLM in cloud deployment scenarios.
+(ii) Under fixed latency constraints, PEVLM increases accuracy from 23.26% to 61.03%, showcasing its critical value for latency sensitive applications.
+Observations
 (a) Full-Attention
-(b) Reuse Position
-(c) Keep Full-Attention Position Embedding
-Figure 1: Comparison of Attention Weight Distributions across Video Contexts.
-The top row shows the attention weights for all tokens, while the bottom row shows those for the video tokens only.
+(b) Parallel Encoding (Reusing Position Embedding)
+(c) Parallel Encoding (Sequential Position Embedding)
+Figure 1: We collected the distribution of attention weights under different scenarios and computed their moving averages for better observation.
+The top row shows the attention weights for all tokens, while the bottom row shows those for the video tokens.
 (a) Full-Attention: The textual tokens and the initial video frames receive the highest attention weights, whereas the remaining video frames are assigned relatively lower attention weights.
-(b) Parallel Encoding with Reused Position Embeddings: The overall attention weights distribution resembles Full-Attention, with higher attention weights for the textual tokens and initial video frames.
+(b) Parallel Encoding (Reusing Position Embedding): The overall attention weights distribution resembles Full-Attention, with higher attention weights for the textual tokens and initial video frames.
 While compared to Full-Attention, the initial video frames receive even higher attention weights, and the remaining video frames exhibit more uniform and lower attention weights.
-(c) Parallel Encoding with Full-Attention Position Embedding: The attention pattern remains similar to (a) and (b), highlighting the textual tokens and the initial video frames.
-However, the attention weight distribution within the video frames more closely aligns with that of Full-Attention.
-Therefore, as illustrated in Figure 1, we perform an attention weight analysis for VLMs and make two key findings: (i) Initial segments of both text and video receive high attention weights, emphasizing the need to include early video frames when selecting sink tokens.
-(ii) Using Full-Attention position embeddings (rather than reusing position embeddings across blocks) yields attention distributions more similar to Full-Attention, leading to better alignment with the model’s intended behavior.
-These insights guide the design of VLM-specific adaptations of parallel-encoding that preserve both efficiency and contextual integrity.
-2.2 Our Contributions
-Basing on these observations, we have conducted a series of studies and experiments.
-And proposed a novel attention mechanism for VLMs named Parallel Encoding for Vision-Language Models (PEVLM), which is a finetune-free method specifically designed to speed up prefill of VLM-based long-video processing.
-By alleviating computational bottlenecks in long video prefill, PEVLM facilitates the efficient deployment of VLMs in various practical application scenarios on both clouds and terminal equipments.
-Our contributions involve:
-  • We introduce PEVLM, a blockwise parallelization strategy for VLMs.
-PEVLM reduces the computational complexity of the attention mechanism from $O((T\times N)^{2})$ to $O(T\times N)$, which leading to a significant improvement in prefill efficiency.
-  • We recover the accuracy of PEVLM with three alignment steps: (i) Segmenting contexts into blocks by video frames rather than tokens.
-(ii) Using the system promots and the initial frames of the video as Sink Block for all blocks to avoid the duplication of abnormal distribution of initial tokens.
-(iii) Preserving the position embeddings of Full-Attention instead of resuing position embeddings across blocks.
-This suggests that PEVLM serves as a performance optimization mechanism rather than endowing the model with the capability to process context lengths that exceed its inherent capacity.
-  • We empirically show that (i) PEVLM gets better accuracy than Block-Attention, APE and Star Attention in long video processing tasks, achieving accuracy improvements ranging from 0.68% to 8.37% across different models.
-(ii) For 100k token contexts, the attention layer achieves a 7.47× speedup, while the end-to-end processing time is reduced by 40% without compromising generation quality.
-This highlights the practical viability and superiority of PEVLM in cloud deployment scenarios.
-(iii) Under fixed latency constraints, PEVLM increases accuracy from 23.26% to 61.03%, showcasing its critical value for latency sensitive applications (e.g., autonomous driving).
-3 PEVLM
-PEVLM is an advanced attention mechanism designed to improve computational efficiency of prefill while maintaining global contextual consistency.
-This approach draws from multiple existing techniques to address computational bottlenecks commonly encountered in long video processing.
-3.1 Standard Attention Mechanism
+(c) Parallel Encoding (Sequential Position Embedding): The overall attention weights distribution remains similar to (a) and (b).
+However, the attention weight distribution within the video frames more closely aligns with (a)).
+To explore the application of parallel encoding in VLMs, we analyzed the potential differences between LLMs and VLMs that may affect accuracy:
+(i) In VLMs, attention sinks are present not only in the initial system prompt, but also in the early video frames  (Kang et al.
+2025).
+To achieve optimal performance, it may be necessary to include enough frames in the sink.
+(ii) Existing parallel-encoding-based methods are primarily designed for Retrieval-Augmented Generation (RAG) tasks (Ma, Wang, and Tian 2025; Yang, Chen, and Chen 2025; Lu et al.
+2024).
+In such tasks, reusing the KV cache from previously retrieved passages is crucial for enhancing inference efficiency.
+Since the order of passages often changes when reusing the KV cache, a common strategy is to apply parallel encoding to generate the KV cache for each passage independently, while sharing position embeddings across passages, as illustrated in Figure 3.
+As a result, position embeddings are typically shared across blocks.
+However, preserving the correct temporal order of video frames is critical for video understanding tasks.
+Blindly reusing position embeddings across blocks can disrupt the model’s ability to capture temporal dependencies.
+This disruption is especially severe for models where the temporal component of the position embeddings is explicitly aligned with absolute time (Bai et al.
+2025).
+Therefore, as illustrated in Figure 1, we perform an attention weight analysis for VLMs and make two key findings:
+(i) As shown in Figure 1(a), initial segments of both text and video receive high attention weights, emphasizing the need to include early video frames when selecting sink tokens.
+(ii) As shown in Figure 1(b), compared to Full-Attention, the initial video frames receive even higher attention weights, while the remaining frames exhibit more uniform and lower attention weights.
+In contrast, when sequential position embeddings are preserved, as illustrated in Figure 1(c), parallel encoding yields an attention score distribution that more closely resembles that of Full-Attention.
+These observations yield two key insights for applying parallel encoding to VLMs: The shared sink block in parallel encoding should include early video frames in addition to textual prompts, as they function as important attention sinks.
+Maintaining the sequential position embeddings of video tokens is essential for preserving temporal consistency; reusing position embeddings across blocks can disrupt temporal alignment.
+These findings underscore the need for a VLM-specific parallel encoding strategy, which we introduce in the next section.
+PEVLM
+With the observation in the last section, we design PEVLM to improve computational efficiency.
+PEVLM enables a seamless shift to parallel encoding without requiring training while maintaining most of the model’s capabilities.
+Our approach adaptively aligns the distribution of attention weights between Full-Attention and parallel encoding, thereby boosting efficiency and performance.
+Standard Attention Mechanism
 In a standard Softmax attention, we attend the query to all past KV states using the following formula:
@@
 Attention Sink.
-First discovered by StreamingLLM (mit2023streamingllm, ), that one or more tokens at the start position of an LLM receive higher attention scores.
-Some VLM-related studies also show that the attention sink phenomenon exists in VLMs but is more distributed among image tokens (kang2025toldvisualattentionsink, ).
+First discovered by StreamingLLM (Lab and AI 2023), that one or more tokens at the start position of an LLM receive higher attention scores.
+APE (Yang, Chen, and Chen 2025) and Star-Attention (Liu et al.
+2024) have demonstrated that sharing a common attention sink across parallel encoding blocks significantly enhances model accuracy.
+Some VLM-related studies also show that the attention sink phenomenon exists in VLMs but is more distributed among image tokens (Kang et al.
+2025).
 Position Embedding.
-To effectively process sequential input, LLMs require position embeddings, such as absolute position embeddings (devlin-etal-2019-bert, )(vaswani2017attention, ) and relative position embeddings (press2022trainshorttestlong, )(SU2024127063, ).
+To effectively process sequential input, LLMs require position embeddings, such as absolute position embeddings (Devlin et al.
+2019; Vaswani et al.
+2017) and relative position embeddings (Press, Smith, and Lewis 2022; Su et al.
+2024).
 For VLMs, contextual positional information is more critical.
 In particular, for some VLM models like Qwen2.5-VL, their position embeddings also incorporate spatio-temporal information between video frames, making position embeddings increasingly essential for VLMs.
-Computational Complexity. In standard Softmax attention mechanisms, the total OPs is
-$OPs_{Attn}=2HL^{2},$ (2)
-where $L$ denotes the context length and $H$ is the hidden size.
-So, the computational complexity is $O(L^{2})$.
-As $L$ increases, the computational cost increases significantly, restricting the inference speed and scalability.
-It is even worse for VLMs.
-Given a video with $T$ frames, where each frame is represented by $N$ tokens (typically, $N$ ranges between 200 and 600), the resulting context length $L$ scales to $(T\times N)$ and complexity to $O((T\times N)^{2})$, making standard attention mechanisms impractical for long video understanding tasks.
-3.2 Parallel Encoding
-3.2.1 Partitioning Strategy
 Figure 2: The PEVLM Masks.
-APE (yang2025ape, ) and Star-Attention (liu2024star, ) have demonstrated that sharing a common attention sink across blocks significantly enhances model accuracy.
-Building on recent advances in efficient attention mechanisms and long-context modeling (mit2023streamingllm, ; liu2024star, ; xiao2023efficient, ; yang2025ape, ), we propose a Sink Block mechanism to improve inference quality.
+Computational Complexity. In standard softmax attention mechanisms, the total number of operations is given by
+$OP_{Attn}=2HL^{2},$ (2)
+where $L$ is the context length and $H$ is the hidden size.
+This leads to a computational complexity of $O(L^{2})$.
+As $L$ increases, the cost grows quadratically, severely limiting the speed and scalability of the inference.
+This issue is further exacerbated in VLMs.
+For a video consisting of $T$ frames, where each frame is represented by $N$ tokens (typically ranging from hundreds to thousands), the total context length becomes $L=T\times N$, resulting in a complexity of $O((T\times N)^{2})$.
+Such scaling renders standard attention mechanisms impractical for long-video understanding tasks.
+Parallel Encoding
+Partitioning Strategy
 As illustrated in Figure 2, our partitioning strategy consists of three core components:
@@
 So we divide the video by frame.
-In addition, as analyzed in Section 2.1, the number of frames selected for the Sink Block depends primarily on the distribution of attention weights.
-We conduct further empirical studies and analyses on this in Section 5.2.
-On the other hand, the number of frames per Context Block involves a trade-off between accuracy and computational efficiency, which we investigate and analyze in Section 5.3.
-3.2.2 Formulations
+Sink Size & Block Size
+In addition, as analyzed in the last section, the number of frames selected for the Sink Block depends primarily on the distribution of attention weights.
+The optimal strategy is to include all tokens with significantly higher attention scores at the beginning of the video in the Sink Block.
+However, as shown in Equation 11, a larger sink size results in higher latency, leading to a trade-off between accuracy and efficiency.
+A similar trade-off also arises in the selection of block size.
+Formulations
 For different blocks, PEVLM applies distinct computation methods to ensure efficient processing and information integration:
@@
 $f(\cdot)$ represents the standard Softmax attention in Equation (1), and $Q$, $K$ and $V$ denote the query, key, and value matrices.
-The total OPs of PEVLM is
-$OPs_{PEVLMAttn}=OPs_{SinkAttn}+N\times OPs_{BlockAttn}+OPs_{QuestAttn},$ (6)
-$OPs_{SinkAttn}=2HS^{2},$ (7)
-$OPs_{BlocksAttn}=2HB(S+B),$ (8)
-$OPs_{QuestAttn}=2HQL,$ (9)
+The total OP of PEVLM is
+$OP_{\text{PEVLM}}=OP_{\text{Sink}}+N\times OP_{\text{ContextBlock}}+OP_{\text{Quest}},$ (6)
+$OP_{\text{Sink}}=2HS^{2},$ (7)
+$OP_{\text{ContextBlock}}=2HB(S+B),$ (8)
+$OP_{\text{Quest}}=2HQL,$ (9)
 where $S$ denotes the sink size, $N$ denotes the block number, $B$ denotes the block size, $Q$ denotes the query size and $H$ is the hidden size. $L$ is the total token number and $L=S+N\times B+Q$.
 After simplification:
-$OPs_{PEVLMAttn}=2H(S^{2}+Q^{2}+NB^{2}+QS+NQB+NSB).$ (10)
-As B is a static number, the computational complexity of PEVLM is $O(L)$.
-3.2.3 Position Encoding in VLMs
+$OP_{\text{PEVLM}}=2H($ $S^{2}+Q^{2}+NB^{2}$ (10)
+$+QS+NQB+NSB).$ (11)
+As the block size is a static number, the computational complexity of PEVLM is $O(L)$.
+Table 1: Performance ($\uparrow$) of different models and different methods on video understanding tasks evaluated at tokens from 26k to 100k.
+Model Method MVBench EgoSchema VideoMME LongVideoBench
+$17s\ in\ avg.$ 3 mins in avg. $<2mins$ $[2mins,1h]$ $<1min$ $[1min,1h]$
+Qwen2.5-VL- Full Attn 68.47% 57.00% 74.56% 56.72% 72.85% 55.33%
+7B-Instruct Block Attn 66.03% 47.40% 65.22% 50.39% 70.64% 47.85%
+Max pixels: 90M APE (T=1.0) 66.03% 57.20% 50.56% 1.06% 54.02% 3.38%
+Tokens: 100k Star Attn 60.19% 19.80% 3.00% 0.00% 49.03% 0.00%
+PEVLM 68.44% 61.00% 74.33% 58.83% 74.52% 57.58%
+LongVILA- Full Attn 61.92% 57.00% 66.33% 49.39% 65.10% 47.44%
+7B-256f Block Attn 57.81% 59.00% 64.89% 51.44% 61.77% 48.05%
+Frames: 256 APE (T=1.0) 60.00% 60.60% 68.56% 51.61% 63.71% 46.93%
+Tokens: 66k Star Attn 63.39% 61.60% 70.11% 54.17% 63.43% 47.44%
+PEVLM 63.53% 61.40% 69.44% 54.33% 65.93% 49.18%
+LLaVA-Video- Full Attn 60.67% 58.20% 75.33% 59.99% 72.02% 55.53%
+7B-Qwen2 (128) Block Attn 56.86% 51.60% 67.56% 55.50% 63.99% 52.36%
+Frames: 128 APE (T=1.0) 58.00% 57.60% 72.22% 58.19% 67.87% 54.61%
+Tokens: 26k Star Attn 59.31% 57.80% 72.89% 58.71% 68.98% 54.92%
+PEVLM 60.00% 58.20% 74.33% 59.99% 72.02% 55.53%
+Position Encoding
 Figure 3: Position Encoding in PEVLM
-As described in Section 2.1, similar to the case in LLMs, applying parallel-encoding-based methods in VLMs suffers from the misalignment of attention weight distributions when compared to Full-Attention.
+As described in observations, similar to the case in LLMs, applying parallel-encoding-based methods in VLMs suffers from the misalignment of attention weight distributions when compared to Full-Attention.
 However, introducing additional hyperparameters to adjust attention weights for better alignment is not ideal for efficient deployment.
-Therefore, inspired by Block Attention (ma2025blockattentionefficientprefilling, ), we retain Full-Attention Position Embedding as illustrated in Figure 3.
-Due to maintaining the original position embedding, PEVLM cannot enable the model to handle longer video contexts beyond its inherent capabilities like some other parallel-encoding-based methods.
-For example, LongVILA-7B-256f can originally process 256 frames at most.
-Although using PEVLM can accelerate the inference speed, it still processes 256 frames at most.
-This is not a problem as our goal is to accelerate the prefill.
-4 Experiments
+Therefore, inspired by Block Attention (Ma, Wang, and Tian 2025) and TurboRAG (Lu et al.
+2024), we retain sequential position embeddings as illustrated in Figure 3.
+Due to retaining the sequential position embeddings, PEVLM does not extend the model’s inherent capacity to handle longer video contexts, unlike some other parallel encoding-based approaches.
+For instance, LongVILA-7B-256f is generally trained to process up to 256 frames, and feeding longer videos may lead to accuracy degradation.
+Although PEVLM accelerates inference, it still processes 256 frames, and longer inputs may still result in degraded performance.
+However, this is acceptable, as our primary objective is to improve prefilling efficiency.
+Experiments
 In this section, we test PEVLM from multiple perspectives against several different VLMs.
 The primary objective is to evaluate the accuracy, computational efficiency, and real-world applicability of PEVLM for long-video processing.
-4.1 Accuracy Evaluation
+And to better reflect practical deployment scenarios, we designed the experiments from two perspectives.
+(i) We first test with workloads of approximately 100k tokens to assess the acceleration performance of PEVLM in cloud serving platforms.
+(ii) Furthermore, we evaluate PEVLM on the LongVideoBench dataset under a given latency constraint, in order to measure its effectiveness in edge scenarios where both computational resources and latency are limited.
+Accuracy Evaluation
+Setup
 The main objective of this experiment is to analyze the impact of PEVLM on accuracy in long-video understanding tasks.
-4.1.1 Setup
-We have chosen the LongVideoBench (longvideobench2024, ) as a benchmark, which is designed to comprehensively evaluate large multimodal models on long-term multimodal understanding.
-We have carried out experiments using LLaVA-Video (zhang2024videoinstructiontuningsynthetic, ), Qwen2.5-VL (qwen2.5-VL, ), and LongVILA (chen2024longvilascalinglongcontextvisual, ) as test models.
-To minimize any adverse effects on model accuracy, all evaluations were performed using bf16 precision.
-As shown in Table 1, we compared the accuracy of Full-Attention, Block-Attention, APE, Star-Attention, and PEVLM in the LongVideoBench dataset.
+We have chosen the LongVideoBench (Zhang and Wang 2024), VideoMME (Fu et al.
+2024), EgoSchema (Mangalam, Akshulakov, and Malik 2023), and MVBench (Li et al.
+2024a) as benchmarks, which are designed to comprehensively evaluate multimodal models in video understanding.
+We have carried out experiments using LLaVA-Video (Zhang et al.
+2024), LongVILA (Chen et al.
+2024), and Qwen2.5-VL (Bai et al.
+2025) as test models.
+To minimize any adverse effects on accuracy, all evaluations were performed using bf16 precision.
+As shown in Table 1, we compared the accuracy of Full-Attention, Block-Attention, APE, Star-Attention, and PEVLM in the datasets.
 Since the other methods are finetune-free, we did not finetune the weights in the Block Attention test.
-For APE and Star-Attention, we tested two scenarios: (i) reusing position encoding across all blocks; (ii) retaining the position encoding of Full-Attention.
+For APE, to accommodate the needs of large-scale deployment, we tested with T=0.
+Regarding the sink block (shared prefix), we selected the system prompt by referring to the code provided by APE.
+For Star-Attention, we implemented its equivalent algorithm on a single node, with the anchor size set equal to the block size, as the configuration in the Block-Attention paper (Ma, Wang, and Tian 2025).
+The block size was uniformly set to 4096 across all algorithms.
+Smaller block sizes resulted in reduced accuracy for all methods, while larger blocks significantly degraded inference performance.
 For PEVLM, we set the sink size and block size to about 16 frames, which corresponds to approximately 4k tokens for Qwen2.5-VL and LongVILA, and about 3k tokens for LLaVA-Video.
-4.1.2 Results
-Table 1: LongVideoBench Accuracy
-model method sink(anchor)_size block_size Correct $\uparrow$
-Full Attn - - 60.43%
-Block Attn (wo finetune) 0 4096 54.23%
-APE 256 4096 NaN
-Qwen2.5-VL- APE (retain pos) 256 4096 60.21%
-7B-Instruct Star Attn 4096 4096 NaN
-Star Attn (retain pos) 4096 4096 61.41%
-PEVLM sys+16frames 16frames 62.23%
-Full Attn - - 52.21%
-Block Attn (wo finetune) 0 4096 51.76%
-APE 256 4096 51.68%
-LongVILA- APE (retain pos) 256 4096 53.18%
-7B-256f Star Attn 4096 4096 46.75%
-Star Attn (retain pos) 4096 4096 53.63%
-PEVLM sys+16frames 16frames 53.70%
-Full Attn - - 59.99%
-Block Attn (wo finetune) 0 4096 55.50%
-APE 256 4096 58.41%
-LLaVA-Video- APE (retain pos) 256 4096 58.49%
-7B-Qwen2 (128) Star Attn 4096 4096 58.71%
-Star Attn (retain pos) 4096 4096 58.49%
-PEVLM sys+16frames 16frames 59.01%
-The experiments demonstrate that PEVLM achieves an accuracy comparable to or even exceeding that of Full-Attention across the three tested models.
-Both APE and Star-Attention can also approach or surpass Full-Attention’s accuracy when retaining Full-Attention’s position encoding.
-However, for Block-Attention, due to the lack of resources for alignment fine-tuning, its accuracy shows a significant decline compared to Full-Attention.
-Without keeping the position encoding of Full-Attention, the output of the Qwen2.5-VL-7B-Instruct model exhibited severe output garbling and semantic confusion, and the LongVILA model also experienced noticeable performance degradation.
-However, the LLaVA-Video model showed insensitivity to changes in position encoding, with negligible accuracy variation observed regardless of whether Full-Attention position encoding was retained.
-4.2 Performance Evaluation
+Results
+The experiments show that PEVLM achieves an accuracy comparable to, and in some cases even surpassing, that of Full-Attention across all three evaluated models.
+Both APE and Star-Attention also approach or exceed the accuracy of Full-Attention when Full-Attention’s position encoding is retained.
+However, Block-Attention exhibits a notable accuracy drop compared to Full-Attention in most settings, primarily due to the lack of alignment-specific fine-tuning.
+Although parallel encoding is generally expected to degrade performance, it can instead improve accuracy in long-context scenarios.
+This is because parallel encoding effectively shortens the total token length involved in each softmax computation, thereby mitigating the softmax feature degradation that occurs with long contexts (Veličković et al.
+2025).
+As a result, the softmax output more robustly approximates an ideal sharp function, which ultimately enhances the performance of some models under long-context conditions.
+Performance Evaluation
 In this section, we evaluate the performance of PEVLM.
-To facilitate experimentation, we built PEVLM on the SGLang framework, which is a fast serving framework for large language models and vision-language models, widely used in cloud production deployment.
-And to better reflect practical deployment scenarios, we designed the experiments from two perspectives.
-(i) We first test with workloads of approximately 100k tokens to assess the acceleration performance of PEVLM in cloud computing environments.
-(ii) Furthermore, we evaluate PEVLM on the LongVideoBench dataset under a given latency constraint, in order to measure its effectiveness in edge scenarios where both computational resources and latency are limited.
-4.2.1 Setup
+To facilitate experimentation, we built PEVLM basied on the SGLang, which is a fast serving framework for large language models and vision-language models, widely used in cloud production deployment.
+Figure 4: We evaluate PEVLM under different block_size settings.
+The sink_size is set to (system prompt+one frame).
+"Attn Time" refers to attention computation time, and "Other Time" covers all other LLM costs.
+"Attn SpeedUP" and "LLM SpeedUP" indicate the acceleration over Attention layer and overall LLM performance.
+Setup
 Since the LLM component in VLM models remains largely consistent across different implementations, performance differences arise primarily from hyperparameter configurations rather than model weights.
 Consequently, we conducted computational efficiency evaluations on the Qwen2.5-VL model.
-The primary objective of this experiment is to assess the improvements in computational efficiency introduced by PEVLM in computational efficiency.
-All experiments were carried out on an H20-96G machine.
+The primary objective of this experiment is to assess the improvements in computational efficiency introduced by PEVLM.
+All experiments were carried out on an NVIDIA H20-96G GPGPU.
 Given that PEVLM primarily optimizes the attention mechanism within the LLM, we measured the execution time of both the entire LLM module and its individual attention layers separately.
-Figure 4: We evaluated the performance of PEVLM under different configurations of (sink_size / block_size).
-Using the Qwen2.5-VL-7B-Instruct model, we tested its LLM and Attention time on a video dataset with an average duration of 10 minutes (approximately 106k tokens per sample).
-"Attn Time" designates the Attention Computation time, whereas "Other Time" signifies the aggregate of all remaining computational costs within the LLM excluding Attention Computation.
-The "Attn SpeedUP" represents the performance acceleration of PEVLM relative to the "Full Attn".
-The "LLM SpeedUP" denotes the overall performance acceleration of the LLM in PEVLM compared to the "Full Attn".
-4.2.2 Results
-As illustrated in Figure 4, the end-to-end inference speed of the LLM is improved by 2.58× to 3.14×.
+Figure 5: We evaluate PEVLM under different sink_size settings.
+The block_size is set to (16 frames).
+"Attn Time" refers to attention computation time, and "Other Time" covers all other LLM costs.
+"Attn SpeedUP" and "LLM SpeedUP" indicate the acceleration over Attention layer and overall LLM performance.
+Results
+As illustrated in Figure 4 and Figure 5, the end-to-end inference speed of the LLM is improved by 2.58× to 3.14×.
 Notably, the attention module achieves a speedup ranging from 7.79× to 33.81×, while the execution time of non-attention components remains largely unaffected.
 As the block size or sink size increases, the acceleration effect of the attention module decreases significantly.
-This observation is consistent with the analysis presented in the Appendix.
+This observation is consistent with the Equation 11.
 Moreover, because non-attention components constitute a substantial portion of the total runtime in PEVLM, the overall end-to-end speedup of the LLM moderately decreases from 3.14× to 2.58×.
-4.3 Accuracy with Limited Latency
+Accuracy with Limited Latency
 One key challenge in deploying VLMs on edge devices is the limited computational resources and strict latency requirements.
@@
 A simplified experimental setup was implemented to assess the efficiency of PEVLM in edge deployment settings.
-4.3.1 Setup
+Table 2: Accuracy with Limited Latency
+Full Attn Full Attn PEVLM
+sink - - sys+2f sys+16f
+block - - 16f 16f
+No Limit 60.43% 55.05% 61.18% 62.23%
+40s 59.69% 55.05% 61.18% 62.23%
+30s 27.08% 55.05% 61.18% 62.23%
+20s 23.26% 24.38% 61.03% 60.28%
+Setup
 We use LongVideoBench as the evaluation benchmark and introduce a latency constraint during inference: any sample whose inference time exceeds the predefined latency threshold is considered a failure.
 The experiments are conducted on the H20-96B platform.
-We select Qwen2.5-VL-7B-Instruct with Full-Attention as the baseline, and compare several methods, including Full-Attention, Block Attention, APE, Star Attention, PEVLM, and using smaller models (Qwen2.5-VL-3B-Instruct).
-Since the other methods are finetune-free, we did not finetune the weights in the Block Attention test.
-Although position encoding does not affect inference lantency, reusing position encoding performs extremely poorly accuracy on the Qwen2.5-VL model.
-Therefore, in the APE and Star Attention tests, we adopted the method of retaining Full-Attention’s position encoding.
-4.3.2 Results
-Table 2: LongVideoBench Accuracy with Limited Latency
-method sink(anchor) block No Limit $\uparrow$ 40s $\uparrow$ 30s $\uparrow$ 20s $\uparrow$
-Full Attn - - 60.43% 59.69% 27.08% 23.26%
-Full Attn (3B) - - 55.05% 55.05% 55.05% 24.38%
-Block Attn 0 4096 54.00% 54.00% 53.78% 53.55%
-APE 256 4096 59.91% 59.91% 59.91% 59.61%
-Star Attn 4096 4096 61.41% 61.41% 61.41% 28.20%
-PEVLM sys+2frames 16frames 61.18% 61.18% 61.18% 61.03%
-PEVLM sys+16frames 16frames 62.23% 62.23% 62.23% 60.28%
+We select Qwen2.5-VL-7B-Instruct with Full-Attention as the baseline, and compare Full-Attention, PEVLM, and using smaller models (Qwen2.5-VL-3B-Instruct).
+Results
 As shown in Table 2, the accuracy of Full-Attention drops significantly when the latency constraint is reduced from 40s to 30s, while the accuracy of Full-Attention 3B also declines sharply when the latency constraint is tightened from 30s to 20s.
-Due to its larger sink and block sizes, StarAttn gets more OPs and exhibits a noticeable accuracy drop between 30s and 20s.
-In contrast, Block Attention, APE, and PEVLM maintain relatively high accuracy even at 20s.
+PEVLM maintain relatively high accuracy even at 20s.
 Notably, although larger sink and block sizes (e.g., sys+16frames, block=16frames) achieve higher accuracy under looser latency constraints, the smaller sink size (sink=sys+2frames, block=16frames) delivers better accuracy at the 20s latency limit due to its lower latency.
-5 How does each component in PEVLM contribute to the performance?
-5.1 Split The Video by Frames instead of Tokens
-Due to the inherent nature of videos being composed of multiple frames, where intra-frame information captures spatial details at a single time point, and inter-frame information represents temporal dynamics, ignoring this structure and arbitrarily dividing blocks based on token count may compromise the spatial integrity at critical temporal boundaries.
-Therefore, we choose to divide the video into blocks based on frames.
-This design is supported by the comparison between Star Attention with Full-Attention positional encoding and PEVLM, as shown in Table 1.
-Across different models, frame-based block partitioning consistently provides accuracy gains, albeit varying in magnitude, over token-based partitioning.
-5.2 Sink Size
-As analyzed in Section 3.2.2, the inference latency of PEVLM increases quadratically with the sink size, which is consistent with the experimental results shown in Figure 4.
+How does each component in PEVLM contribute to the performance?
+In Table 3, we present an ablation study to evaluate the contribution of each component in PEVLM, including the use of Sequential Position Embedding (SP), Dividing the Video by Frames (DF), and Adding Frames to the Sink (FS).
+The results are averaged across the four base models reported in Table 1.
+We observe that sequential parallel encoding yields the most significant improvement in accuracy, achieving an average performance gain of 12.64%.
+Incorporating the initial frames into the sink block leads to an additional average accuracy gain of 1.54%.
+In contrast, dividing the video by frames provides only a marginal improvement compared to directly segmenting the video without this strategy.
+Table 3: Ablation study of PEVLM components on LongVideoBench
+SP DF FS LongVideoBench VideoMME
+46.67% 45.31%
+$\checkmark$ 56.75% 60.51%
+$\checkmark$ $\checkmark$ 56.77% 60.63%
+$\checkmark$ $\checkmark$ $\checkmark$ 58.26% 62.22%
+Position Encoding
+We further analyze the component with the greatest impact on accuracy: the use of sequential position embeddings.
+As discussed before, adopting sequential position embeddings instead of reusing position embeddings across blocks produces attention distributions that more closely resemble those of Full-Attention, thereby aligning better with the model’s expected behavior.
+As shown in Table 4, the use of sequential position embeddings significantly improves accuracy in both Qwen2.5-VL and LongVILA.
+In contrast, for the LLaVA-Video model, the benefit is marginal, likely because its attention weight distribution (as shown in Figure 9 in the Appendix) remains relatively stable even when using parallel encoding.
+Table 4: Accuracy $\uparrow$ on LongVideoBench dataset with Different Position Embedding Strategy
+model method reuse pos sequential
+Qwen2.5-VL- APE (T=1.0) 30.44% 59.39%
+7B-Instruct Star Attn 29.54% 61.41%
+PEVLM 14.73% 62.23%
+APE (T=1.0) 51.46% 53.40%
+7B-256f Star Attn 51.76% 53.63%
+PEVLM 49.96% 53.70%
+LLaVA-Video- APE (T=1.0) 58.19% 58.12%
+7B-Qwen2 Star Attn 58.71% 58.49%
+(128) PEVLM 58.34% 59.01%
+Sink & Block Size
+As analyzed in the third section, the inference latency of PEVLM increases quadratically with the sink size, which is consistent with the experimental results shown in Figure 5.
 Furthermore, we investigate the impact of sink size on model accuracy.
-As observed in Section 2.1, it is essential to involve the initial frames of the video in the sink.
+As observed in the second section, it is essential to involve the initial frames of the video in the sink.
 This observation is validated in Figure 6, where incorporating the initial frames into the sink leads to varying degrees of accuracy improvement across different models.
-Moreover, as shown in Figure 5, the accuracy of Qwen2.5-VL and LLaVA-Video shows an upward trend as the sink size increases, though the improvement remains marginal.
-In contrast, LongVILA exhibits fluctuations in accuracy with varying sink sizes but does not show a consistent upward or downward trend.
+Figure 6: The impact of sink sizes on accuracy
+Moreover, as shown in Figure 6, the accuracy of all models shows an upward trend as the sink size increases.
 However, because of the quadratic increase in latency, using longer sink sizes is not favorable for performance.
@@
 If latency constraints are also taken into account, this could lead to a trade-off between performance and accuracy.
-For example, as demonstrated in the results of Section 4.3, under a 20-second latency constraint, using only sys+2frames as the Sink Block achieves better accuracy than using sys+16frames.
-Figure 5: The impact of sink/block sizes on accuracy
-5.3 Block Size
-As analyzed in Section 3.2.2, the inference latency of PEVLM also increases quadratically with the block size as well as the sink size, which is consistent with the experimental results shown in Figure 4.
-However, the block size has a more significant impact on performance because the coefficient of block-size’s quadratic term in Equation 10 is much larger than that of the sink size.
+For example, as demonstrated in the results of Table 2, under a 20-second latency constraint, using only sys+2frames as the Sink Block achieves better accuracy than using sys+16frames.
+Figure 7: The impact of block sizes on accuracy
+As shown in Figure 4, the inference latency of PEVLM also increases quadratically with block size, which is consistent with the experimental results shown in Figure 4.
+However, the block size has a more significant impact on performance because the coefficient of the block size’s quadratic term in Equation 11 is much larger than that of the sink size.
 We investigate the impact of block size on model accuracy here.
@@
 However, due to the quadratic growth in latency, a trade-off arises between computational efficiency and accuracy.
-5.4 Position Encoding
-As analyzed in Section 2.1, using Full-Attention position embeddings, rather than reusing position embeddings across blocks, should yield attention distributions that more closely resemble those of Full-Attention, thereby aligning better with the model’s intended behavior.
-This is validated by the experiments presented in Section 4.1.
-As shown in Table 1, for several parallel-encoding-based methods (Block-Attention, APE, Star Attention, and PEVLM), using Full-Attention position embeddings significantly improves accuracy in both Qwen2.5-VL and LongVILA.
-In contrast, for the LLaVA-Video model, the benefit is marginal, likely because its attention weight distribution remains relatively stable even when using parallel encoding.
-6 Conclusion
+Conclusion
 This study investigates the potential of parallel-encoding-based methods in VLMs, which can reduce the computational complexity of attention mechanisms for fast inference and reuse positions for long-context processing but often result in performance degradation.
 To address this issue, we propose PEVLM, a training-free method designed to enable accurate, fast, and long-context video processing.
-PEVLM achieves this by aligning the attention weight distribution of parallel encoding with Full-Attention through two key steps: selecting an appropriate sink size and block size, and utilizing Full-Attention position embeddings.
+PEVLM achieves this by aligning the attention weight distribution of parallel encoding with Full-Attention through two key steps: selecting an appropriate sink size and block size, and utilizing sequential position embeddings.
 Empirically, we demonstrate that PEVLM enhances both accuracy and efficiency in long video processing tasks while successfully scaling to process hundreds of chunks in parallel across various settings.
-7 Future Work
+Future Work
 Although PEVLM has demonstrated substantial improvements in computational efficiency and reasoning performance for long video processing tasks, there remains ample opportunity for further research and enhancement.
 In future work, we plan to explore the following directions.
-In autonomous driving scenarios, a common strategy is to use several historical frames as input to make decisions for the next action.
-Due to the computational nature of the attention mechanism, most current approaches require re-computation of all historical frames as the temporal step advances.
-Building upon the current PEVLM framework, we propose integrating the rotating positional embedding method from Block-Attention (ma2025blockattentionefficientprefilling, ), which enables the reuse of the KVCache from all historical frames preceding the current timestep.
-This approach has the potential to yield an order-of-magnitude improvement in efficiency for such applications.
-Many of the existing methods based on parallel-encoding, including PEVLM, can achieve accuracy that exceeds Full-Attention.
-This phenomenon is attributed to the fact that parallel-encoding-based methods can mitigate the "lost in the middle" problem, thereby enabling them to achieve accuracy surpassing Full-Attention when processing long contexts (yang2025ape, ).
-The reason why parallel-encoding can mitigate the "lost in the middle" issue is hypothesized to be due to its shortening of the distance between the context and the query.
-However, more concrete evidence is needed to support this claim.
+Beyond video inputs, an important future direction is extending PEVLM to handle richer multimodal inputs such as LiDAR, map priors, and inertial measurements, which are commonly used in domains like autonomous driving and robotics.
+These modalities often exhibit different spatial and temporal characteristics compared to visual tokens, posing unique alignment and fusion challenges for attention-based models.
+Enhancing PEVLM with cross-modal position encoding strategies and adaptive sink block selection could further improve its generalization to complex, heterogeneous input streams.
+One promising direction is extending PEVLM to streaming or online scenarios, such as autonomous driving, where the model continuously receives multimodal inputs (e.g., video frames and LiDAR) over time.
+In these settings, a fixed-length sliding temporal window is often adopted: new data is appended while the oldest is discarded, maintaining a constant context size that reflects the most recent few seconds or minutes of input.
+If PEVLM can be adapted to such settings, it would allow the model to compute only over the newly arrived data at each inference step, rather than reprocessing the entire history.
+This could lead to significant performance gains—potentially improving inference speed by several times or even orders of magnitude.
+Such an extension would enable real-time video-language reasoning under strict latency constraints and dynamic, continuously evolving environments.
+Appendix A Benchmark Details
+We evaluate our method on several video understanding benchmarks that test different aspects of video comprehension:
+EgoSchema
+EgoSchema (Mangalam, Akshulakov, and Malik 2023) is a large-scale benchmark designed to evaluate Multimodal Large Language Models (MLLMs) on egocentric video understanding.
+It consists of 100 hours of first-person video data spanning 1,270 daily activity episodes across diverse real-world environments.
+The benchmark introduces over 10,000 manually curated question-answer pairs, covering tasks such as object grounding, human-object interaction, activity reasoning, and intent prediction.
+MVBench
+MVBench (Li et al.
+2024a) is a comprehensive benchmark designed to evaluate Multimodal Large Language Models (MLLMs) on multi-granular video understanding.
+It consists of 5 task categories—including moment-level, frame-level, clip-level, video-level, and holistic video understanding—covering a wide range of temporal scopes and reasoning demands.
+The benchmark includes 2,562 manually annotated questions grounded in 4,119 diverse video clips, selected from real-world scenarios.
+Each question is carefully designed to probe different levels of spatiotemporal understanding, from fine-grained object recognition and short-term motion tracking to long-term event inference.
+MVBench offers a unified and challenging evaluation protocol to assess the generalization and reasoning ability of MLLMs across granularities.
+Unlike conventional third-person video datasets, EgoSchema emphasizes embodied perception and temporal reasoning from an egocentric perspective, posing unique challenges for spatial understanding, long-term memory, and causal inference in MLLMs.
+Video-MME
+Video-MME (Fu et al.
+2024) is a comprehensive evaluation benchmark for assessing the video understanding capabilities of Multimodal Large Language Models (MLLMs).
+It spans 6 primary visual domains and 30 subfields, covering a diverse range of video types and temporal scenarios.
+The benchmark includes 900 videos with durations ranging from 11 seconds to 1 hour, totaling 254 hours of content.
+To evaluate fine-grained visual and temporal reasoning, 2,700 manually annotated question-answer pairs are provided.
+Video-MME challenges models to comprehend both short and long video clips across different temporal granularities, making it a rigorous benchmark for evaluating the core video processing capabilities of MLLMs.
+LongVideoBench
+LongVideoBench (Zhang and Wang 2024) is a benchmark specifically designed to evaluate the long-context video understanding capabilities of Multimodal Large Language Models (MLLMs).
+It features 1,760 videos spanning 12 diverse real-world scenarios, with video durations ranging from 5 minutes to 2 hours, totaling over 1,000 hours of content.
+The benchmark includes 2,400 manually annotated multi-choice questions targeting key aspects of long video comprehension.
+LongVideoBench focuses on challenging long-range temporal reasoning, event tracking, and global understanding across extended video content.
+It aims to assess whether models can maintain coherence, memory, and attention over prolonged contexts, making it a rigorous testbed for long-form video modeling.
+Appendix B Attention Weights Distribution of LLaVA-Video and LongVILA
+(a) Full-Attention
+(b) Reuse position embeddings
+(c) Sequential position embeddings
+Figure 8: Attention Weight Distributions of LongVILA
+(a) Full-Attention
+(b) Reuse position embeddings
+(c) Sequential position embeddings
+Figure 9: Attention Weight Distributions of LLaVA-Video

### v3

@@
 Introduction
-In recent years, Vision-Language Models (VLMs) have emerged as a central research focus at the intersection of computer vision and natural language processing, demonstrating impressive performance in multimodal understanding and generation tasks (Alayrac et al.
+In recent years, Vision-Language Models (VLMs) have become a central research focus at the intersection of computer vision and natural language processing.
+These models have demonstrated impressive performance across a wide range of multimodal understanding and generation tasks (Alayrac et al.
 2022; Li et al.
@@
 2023).
-As the capabilities of these models continue to strengthen, their application scenarios have also expanded, including robotics (Black, Lee et al.
+As their capabilities continue to improve, VLMs are being applied in increasingly complex domains, including robotics (Black, Lee et al.
 2024; Cheang, Huang et al.
@@
 2024).
-Correspondingly, the length of the videos has also increased significantly.
-However, applying VLMs to long-video data introduces a significant computational challenge: The quadratic complexity of transformer-based attention mechanisms causes computation to scale rapidly with the input sequence length (Beltagy, Peters, and Cohan 2020; Vaswani et al.
-2017), especially during the prefilling stage, where all input tokens must be processed before generating any output.
-This severely limits the scalability of VLMs in long-video understanding tasks.
-In practical scenarios where the entire video must be encoded before decoding begins—such as video captioning, retrieval, or reasoning—this inefficiency becomes a critical bottleneck.
-Therefore, optimizing the prefilling phase is essential to efficiently deploy VLMs on long video tasks.
-A widely adopted technique in LLMs to speed up long context prefilling is the parallel encoding mechanism (Li et al.
+These application scenarios often demand processing longer video inputs, placing significant strain on computational resources.
+A key obstacle to applying VLMs to long-video inputs is the quadratic complexity of transformer attention during the prefilling stage (Vaswani et al.
+2017; Beltagy, Peters, and Cohan 2020).
+This computational burden limits scalability in tasks such as perception, temporal reasoning, and action planning, which are fundamental to domains like robotics, autonomous driving, and healthcare.
+As a result, optimizing the prefilling phase is critical to enabling efficient long-video processing with VLMs.
+To address this issue, a widely adopted technique in large language models (LLMs) is the parallel encoding mechanism (Li et al.
 2024b; Liu et al.
@@
 2024).
-In these approaches, the input context is divided into blocks, and each block is encoded into Key-Value (KV) states independently.
-This means that tokens from different blocks do not attend to each other.
-Since the parallel encoding mechanism omits attention computation across blocks, it reduces the computational complexity from $O(L^{2})$ to $O(L)$.
-The parallel encoding mechanism also mitigates the problem of "lost in the middle" (Liu et al.
-2023), and in long-context scenarios, it can even achieve accuracy surpassing that of Full-Attention (Yang, Chen, and Chen 2025).
+In this approach, the input context is divided into multiple blocks, with each block independently encoded into Key-Value (KV) states.
+This design prevents tokens in different blocks from attending to one another, thereby reducing the computational complexity from $O(L^{2})$ to $O(L)$.
+Moreover, parallel encoding alleviates the issue of “lost in the middle” by reducing the number of tokens participating in the softmax operation and can even achieve accuracy surpassing that of Full-attention in long-context scenarios (Liu et al.
+2023; Yang, Chen, and Chen 2025; Veličković et al.
+2025).
 Several state-of-the-art architectural improvements have been proposed to enhance the parallel encoding mechanism.
@@
 The issue is particularly pronounced in domains such as autonomous driving, where a reliable models are often continuously trained and iterated over years.
-Consequently, we focus on finetune-free optimization methods to ensure compatibility and ease of deployment.
+Consequently, we focus on finetune-free performance optimization methods to ensure compatibility and ease of deployment.
 Given all of these, a key question naturally arises: Can existing methods be directly applied to VLMs to accelerate inference and deployment pipelines?
@@
 2024a) benchmarks.
-Although existing parallel encoding methods achieve some effectiveness in shorter videos, we observe varying degrees of accuracy degradation across different models on longer videos, particularly for Qwen2.5-VL.
+Although existing parallel encoding methods demonstrate certain effectiveness on shorter videos, we observe varying degrees of accuracy degradation across different models when applied to longer videos.
+This performance drop is particularly pronounced in models like Qwen2.5-VL, which are more sensitive to spatial and temporal information in videos—sometimes resulting in completely empty or garbled outputs in long-video scenarios, with accuracy dropping to 0.0%.
 To investigate this issue, we analyze the attention weight distributions under parallel encoding and identify a misalignment between Full-Attention and parallel encoding.
 Existing methods developed for LLMs fail to effectively address this misalignment in VLMs.
-Building upon this observation.
-We conducted a series of studies and experiments and ultimately propose a novel attention mechanism for VLMs, termed Parallel Encoding for Vision-Language Models (PEVLM).
+Building upon this observation, we conducted a series of studies and experiments and ultimately propose a novel attention mechanism for VLMs, termed Parallel Encoding for Vision-Language Models (PEVLM).
 PEVLM is a fine-tuning-free method specifically designed to accelerate the prefilling stage in long-video processing with VLMs.
@@
 Our observations reveal a key reason for the misalignment between the attention weight distributions of parallel encoding and Full-Attention: the reuse of position embeddings across blocks leads to the loss of critical video information.
-  • We propose PEVLM to recover the accuracy of parallel encoding by applying three targeted alignment steps: (i) Segmenting contexts into blocks by video frames rather than tokens.
-(ii) Using the system promots and the initial frames of the video as Sink Block for all blocks to avoid the duplication of abnormal distribution of initial tokens.
-(iii) Preserving the sequential position embeddings instead of resuing position embeddings across blocks.
-PEVLM achieves superior accuracy compared to other parallel-encoding-based methods on long video processing tasks, with up to a 9.73% improvement across different models.
+  • We propose PEVLM to recover the accuracy of parallel encoding by applying three targeted alignment steps: (i) segmenting contexts into blocks by video frames rather than tokens; (ii) using the system promots and the initial frames of the video as Sink Block for all blocks to avoid the duplication of abnormal distribution of initial tokens; (iii) preserving the sequential position embeddings instead of resuing position embeddings across blocks.
+With these alignment strategies, PEVLM achieves higher accuracy than existing parallel encoding methods, and reaches 99.57% to 104.80% of the accuracy of full attention on different models.
   • PEVLM reduces the computational complexity of the attention mechanism from $O((T\times N)^{2})$ to $O(T\times N)$, which leading to a significant improvement in prefilling efficiency.
-(i) For 100k token contexts prefilling, the attention layer achieves a 7.47× speedup, while the end-to-end 2.58x speedup without compromising generation quality.
+(i) For 100k (text&video) token contexts prefilling, the attention layer achieves a 7.47× speedup, while the end-to-end 2.58x speedup without compromising generation quality.
 This highlights the practical viability and superiority of PEVLM in cloud deployment scenarios.
@@
 Figure 1: We collected the distribution of attention weights under different scenarios and computed their moving averages for better observation.
-The top row shows the attention weights for all tokens, while the bottom row shows those for the video tokens.
+The top row shows the attention weights for all tokens, while the bottom row shows those for the visual tokens.
 (a) Full-Attention: The textual tokens and the initial video frames receive the highest attention weights, whereas the remaining video frames are assigned relatively lower attention weights.
@@
 Therefore, as illustrated in Figure 1, we perform an attention weight analysis for VLMs and make two key findings:
-(i) As shown in Figure 1(a), initial segments of both text and video receive high attention weights, emphasizing the need to include early video frames when selecting sink tokens.
+(i) As shown in Figure 1(a), some initial text and visual tokens receive high attention scores, consistent with the observations reported in (Lab and AI 2023; Sun et al.
+2024; Kang et al.
+2025).
+This suggests that it may be necessary to include early video frames when selecting sink tokens.
 (ii) As shown in Figure 1(b), compared to Full-Attention, the initial video frames receive even higher attention weights, while the remaining frames exhibit more uniform and lower attention weights.
 In contrast, when sequential position embeddings are preserved, as illustrated in Figure 1(c), parallel encoding yields an attention score distribution that more closely resembles that of Full-Attention.
-These observations yield two key insights for applying parallel encoding to VLMs: The shared sink block in parallel encoding should include early video frames in addition to textual prompts, as they function as important attention sinks.
-Maintaining the sequential position embeddings of video tokens is essential for preserving temporal consistency; reusing position embeddings across blocks can disrupt temporal alignment.
+These observations yield two key insights for applying parallel encoding to VLMs: (i) the shared sink block in parallel encoding should include early video frames in addition to textual prompts, as they function as important attention sinks; (ii) maintaining the sequential position embeddings of visual tokens is essential for preserving temporal consistency, and reusing position embeddings across blocks could disrupt temporal alignment.
 These findings underscore the need for a VLM-specific parallel encoding strategy, which we introduce in the next section.
@@
 With the observation in the last section, we design PEVLM to improve computational efficiency.
-PEVLM enables a seamless shift to parallel encoding without requiring training while maintaining most of the model’s capabilities.
+PEVLM enables a seamless shift from Full-Attention to parallel encoding without requiring training while maintaining most of the model’s capabilities.
 Our approach adaptively aligns the distribution of attention weights between Full-Attention and parallel encoding, thereby boosting efficiency and performance.
 Standard Attention Mechanism
-In a standard Softmax attention, we attend the query to all past KV states using the following formula:
+Before introducing PEVLM, we briefly review key aspects of the standard attention mechanism as a foundation. The computation formula for standard softmax attention is:
 $O=Softmax(\frac{QK^{T}}{\sqrt{d}})V\ \ \ Q\in\mathbb{R}^{n\times d}\ \ \ K,V\in\mathbb{R}^{m\times d}$ (1)
@@
 Attention Sink.
-First discovered by StreamingLLM (Lab and AI 2023), that one or more tokens at the start position of an LLM receive higher attention scores.
+The attention sink phenomenon was first identified by StreamingLLM (Lab and AI 2023), which observed that one or more tokens at the beginning of the input consistently receive higher attention scores.
 APE (Yang, Chen, and Chen 2025) and Star-Attention (Liu et al.
-2024) have demonstrated that sharing a common attention sink across parallel encoding blocks significantly enhances model accuracy.
-Some VLM-related studies also show that the attention sink phenomenon exists in VLMs but is more distributed among image tokens (Kang et al.
+2024) have demonstrated that sharing a common attention sink across parallel encoding blocks significantly enhances the accuracy of the model.
+Some studies also show that the attention sink phenomenon exists in VLMs but is more distributed among visual tokens (Kang et al.
 2025).
+Figure 2: The PEVLM Masks.
 Position Embedding.
-To effectively process sequential input, LLMs require position embeddings, such as absolute position embeddings (Devlin et al.
-2019; Vaswani et al.
-2017) and relative position embeddings (Press, Smith, and Lewis 2022; Su et al.
-2024).
+Position embedding provides essential information that distinguishes the position and order amongst tokens in a sequence.
 For VLMs, contextual positional information is more critical.
-In particular, for some VLM models like Qwen2.5-VL, their position embeddings also incorporate spatio-temporal information between video frames, making position embeddings increasingly essential for VLMs.
-Figure 2: The PEVLM Masks.
-Computational Complexity. In standard softmax attention mechanisms, the total number of operations is given by
+In particular, for some VLM models like Qwen2.5-VL (Bai et al.
+2025), their position embeddings also incorporate spatio-temporal information between video frames, making position embeddings increasingly essential for VLMs.
+Computational Complexity. In standard softmax attention mechanisms, the total number of operations in the prefilling phase is given by
 $OP_{Attn}=2HL^{2},$ (2)
@@
   • Question Block: The text tokens that follow the video input are left unsegmented as Quesion Block.
+Position Encoding
+Figure 3: Position Encoding in PEVLM
+As described in the Observations section, similar to the case in LLMs, applying parallel encoding methods in VLMs suffers from the misalignment of attention weight distributions when compared to Full-Attention.
+While prior work has attempted to address this issue by introducing additional hyperparameters to realign attention scores (Yang, Chen, and Chen 2025), such approaches are not ideal for efficient deployment.
+To address this, PEVLM draws inspiration from Block Attention (Ma, Wang, and Tian 2025) and TurboRAG (Lu et al.
+2024), and retains sequential position embeddings, as shown in Figure 3.
+In contrast to these methods, which dynamically adjust positional embeddings, often to accommodate context reuse through sequence reordering, PEVLM forgoes such updates.
+Since video frames are processed in their original temporal order, PEVLM directly applies the sequential position embedding to each visual token within context blocks, without modifying their positions after query concatenation.
+By preserving sequential position embeddings, PEVLM maintains attention weight distributions that closely resemble those of Full Attention.
+However, unlike some parallel encoding methods that aim to extend context capacity, this design choice also constrains PEVLM’s ability to scale to longer video sequences.
+For instance, LongVILA-7B-256f is typically trained on sequences up to 256 frames; attempting to process longer inputs may lead to performance degradation.
+Although PEVLM achieves significant acceleration of inference, it is similarly limited to 256-frame inputs.
+Nonetheless, this trade-off is acceptable as our primary goal is to enhance inference efficiency rather than to expand context length.
+Table 1: Performance ($\uparrow$) of different models and different methods on video understanding tasks evaluated at tokens from 26k to 100k.
+Model Method MVBench EgoSchema VideoMME LongVideoBench (Val) Avg.
+[…]
+Qwen2.5-VL- Full Attn 68.47% 57.00% 74.56% 56.72% 72.85% 55.33% 64.16%
+7B-Instruct Block Attn 66.03% 47.40% 65.22% 50.39% 70.64% 47.85% 57.92%
+Frames: 768 APE(T=1.0) 66.03% 57.20% 50.56% 1.06% 54.02% 3.38% 38.71%
+Tokens: 100k Star Attn 60.19% 19.80% 3.00% 0.00% 49.03% 0.00% 22.00%
+PEVLM 68.44% 61.00% 74.33% 58.83% 74.52% 57.58% 65.78%
+LongVILA- Full Attn 61.92% 57.00% 66.33% 49.39% 65.10% 47.44% 57.86%
+7B-256f Block Attn 57.81% 59.00% 64.89% 51.44% 61.77% 48.05% 57.16%
+Frames: 256 APE(T=1.0) 60.00% 60.60% 68.56% 51.61% 63.71% 46.93% 58.57%
+Tokens: 66k Star Attn 63.39% 61.60% 70.11% 54.17% 63.43% 47.44% 60.02%
+PEVLM 63.53% 61.40% 69.44% 54.33% 65.93% 49.18% 60.64%
+LLaVA-Video- Full Attn 60.67% 58.20% 75.33% 59.99% 72.02% 55.53% 63.62%
+7B-Qwen2 Block Attn 56.86% 51.60% 67.56% 55.50% 63.99% 52.36% 57.98%
+Frames: 128 APE(T=1.0) 58.00% 57.60% 72.22% 58.19% 67.87% 54.61% 61.42%
+Tokens: 26k Star Attn 59.31% 57.80% 72.89% 58.71% 68.98% 54.92% 62.10%
+PEVLM 60.00% 58.20% 74.33% 59.99% 72.02% 55.53% 63.35%
+Formulations
+The computational formulation of PEVLM is defined as:
+$\text{Attn}s=f(Q_{s},K_{s},V_{s}),$ (3)
+$\text{Attn}_{c_{i}}=f(Q_{c_{i}},K_{s+c_{0}+\dots+c_{i-1}},V_{s+c_{0}+\dots+c_{i-1}}),$ (4)
+$\text{Attn}_{q}\ =f(Q_{q},K_{s+c_{\text{all}}+q},V_{s+c_{\text{all}}+q}),$ (5)
+where $s$ denotes the Sink Block, $q$ denotes the Question Block, and $c_{i}$ denotes the $i$-th Context Block.
+Symbols $Attn_{s}$, $Attn_{c_{i}}$, and $Attn_{q}$ represent the attention outputs for the corresponding blocks.
+The function $f(\cdot)$ corresponds to the standard softmax attention defined in Equation (1), with $Q$, $K$ and $V$ denoting the query, key, and value matrices respectively.
+The total computational operation (OP) count of PEVLM is:
+$\text{OP}_{\text{PEVLM}}=\text{OP}_{\text{Sink}}+N\times\text{OP}_{\text{ContextBlock}}+\text{OP}_{\text{Quest}},$ (6)
+$\text{OP}_{\text{Sink}}=2HS^{2},$ (7)
+$\text{OP}_{\text{ContextBlock}}=2HB(S+B),$ (8)
+$\text{OP}_{\text{Quest}}=2HQL,$ (9)
+where $S$ denotes the sink block size, $N$ denotes the context block number, $B$ denotes the context block size, $Q$ denotes the query block size and $H$ is the hidden size.
+$L$ is the total token number and $L=S+N\times B+Q$.
+After simplification:
+$\text{OP}_{\text{PEVLM}}=2H($ $S^{2}+Q^{2}+NB^{2}$
+$+QS+NQB+NSB).$ (10)
+As the context block size is a static number, the computational complexity of PEVLM is $O(L)$.
+Block Size
 Since the information within each video frame represents spatial content at a specific moment, while information across frames reflects temporal dynamics, videos naturally possess structural boundaries.
@@
 So we divide the video by frame.
-Sink Size & Block Size
-In addition, as analyzed in the last section, the number of frames selected for the Sink Block depends primarily on the distribution of attention weights.
+As analyzed in the previous section, the number of frames selected for the Sink Block depends primarily on the distribution of attention weights.
 The optimal strategy is to include all tokens with significantly higher attention scores at the beginning of the video in the Sink Block.
-However, as shown in Equation 11, a larger sink size results in higher latency, leading to a trade-off between accuracy and efficiency.
-A similar trade-off also arises in the selection of block size.
-Formulations
-For different blocks, PEVLM applies distinct computation methods to ensure efficient processing and information integration:
-$\text{Attn}_{s}\ =f(Q_{s},K_{s},V_{s}),$ (3)
-$\text{Attn}_{c_{i}}=f(Q_{c_{i}},K_{s+c_{0}+\dots+c_{i-1}},V_{s+c_{0}+\dots+c_{i-1}}),$ (4)
-$\text{Attn}_{q}\ =f(Q_{q},K_{s+c_{\text{all}}+q},V_{s+c_{\text{all}}+q}),$ (5)
-where $s$ and $q$ are the Sink Block and question block, $c_{i}$ is the ith Context Blocks; $Attn_{s}$, $Attn_{c_{i}}$ and $Attn_{q}$ denote the attention of the $s$, $c_{i}$ and $q$.
-$f(\cdot)$ represents the standard Softmax attention in Equation (1), and $Q$, $K$ and $V$ denote the query, key, and value matrices.
-The total OP of PEVLM is
-$OP_{\text{PEVLM}}=OP_{\text{Sink}}+N\times OP_{\text{ContextBlock}}+OP_{\text{Quest}},$ (6)
-$OP_{\text{Sink}}=2HS^{2},$ (7)
-$OP_{\text{ContextBlock}}=2HB(S+B),$ (8)
-$OP_{\text{Quest}}=2HQL,$ (9)
-where $S$ denotes the sink size, $N$ denotes the block number, $B$ denotes the block size, $Q$ denotes the query size and $H$ is the hidden size. $L$ is the total token number and $L=S+N\times B+Q$.
-After simplification:
-$OP_{\text{PEVLM}}=2H($ $S^{2}+Q^{2}+NB^{2}$ (10)
-$+QS+NQB+NSB).$ (11)
-As the block size is a static number, the computational complexity of PEVLM is $O(L)$.
-Table 1: Performance ($\uparrow$) of different models and different methods on video understanding tasks evaluated at tokens from 26k to 100k.
-Model Method MVBench EgoSchema VideoMME LongVideoBench
-$17s\ in\ avg.$ 3 mins in avg. $<2mins$ $[2mins,1h]$ $<1min$ $[1min,1h]$
-Qwen2.5-VL- Full Attn 68.47% 57.00% 74.56% 56.72% 72.85% 55.33%
-7B-Instruct Block Attn 66.03% 47.40% 65.22% 50.39% 70.64% 47.85%
-Max pixels: 90M APE (T=1.0) 66.03% 57.20% 50.56% 1.06% 54.02% 3.38%
-Tokens: 100k Star Attn 60.19% 19.80% 3.00% 0.00% 49.03% 0.00%
-PEVLM 68.44% 61.00% 74.33% 58.83% 74.52% 57.58%
-LongVILA- Full Attn 61.92% 57.00% 66.33% 49.39% 65.10% 47.44%
-7B-256f Block Attn 57.81% 59.00% 64.89% 51.44% 61.77% 48.05%
-Frames: 256 APE (T=1.0) 60.00% 60.60% 68.56% 51.61% 63.71% 46.93%
-Tokens: 66k Star Attn 63.39% 61.60% 70.11% 54.17% 63.43% 47.44%
-PEVLM 63.53% 61.40% 69.44% 54.33% 65.93% 49.18%
-LLaVA-Video- Full Attn 60.67% 58.20% 75.33% 59.99% 72.02% 55.53%
-7B-Qwen2 (128) Block Attn 56.86% 51.60% 67.56% 55.50% 63.99% 52.36%
-Frames: 128 APE (T=1.0) 58.00% 57.60% 72.22% 58.19% 67.87% 54.61%
-Tokens: 26k Star Attn 59.31% 57.80% 72.89% 58.71% 68.98% 54.92%
-PEVLM 60.00% 58.20% 74.33% 59.99% 72.02% 55.53%
-Position Encoding
-Figure 3: Position Encoding in PEVLM
-As described in observations, similar to the case in LLMs, applying parallel-encoding-based methods in VLMs suffers from the misalignment of attention weight distributions when compared to Full-Attention.
-However, introducing additional hyperparameters to adjust attention weights for better alignment is not ideal for efficient deployment.
-Therefore, inspired by Block Attention (Ma, Wang, and Tian 2025) and TurboRAG (Lu et al.
-2024), we retain sequential position embeddings as illustrated in Figure 3.
-Due to retaining the sequential position embeddings, PEVLM does not extend the model’s inherent capacity to handle longer video contexts, unlike some other parallel encoding-based approaches.
-For instance, LongVILA-7B-256f is generally trained to process up to 256 frames, and feeding longer videos may lead to accuracy degradation.
-Although PEVLM accelerates inference, it still processes 256 frames, and longer inputs may still result in degraded performance.
-However, this is acceptable, as our primary objective is to improve prefilling efficiency.
+However, on the one hand, the number of frames containing visual tokens with a high attention score varies across different videos; on the other hand, as shown in Equation 10, a larger sink block size leads to higher latency, resulting in a trade-off between accuracy and efficiency.
+A similar trade-off also exists in determining the context block size.
+We will further analyze this in the following sections.
 Experiments
@@
 Accuracy Evaluation
+This experiment aims to evaluate the impact of PEVLM on accuracy in long-video understanding tasks.
 Setup
-The main objective of this experiment is to analyze the impact of PEVLM on accuracy in long-video understanding tasks.
-We have chosen the LongVideoBench (Zhang and Wang 2024), VideoMME (Fu et al.
+We adopt LongVideoBench (Zhang and Wang 2024), VideoMME (Fu et al.
 2024), EgoSchema (Mangalam, Akshulakov, and Malik 2023), and MVBench (Li et al.
 2024a) as benchmarks, which are designed to comprehensively evaluate multimodal models in video understanding.
-We have carried out experiments using LLaVA-Video (Zhang et al.
+All evaluations are conducted using the lmms-eval toolkit (Li* et al.
+2024).
+We test on LLaVA-Video (Zhang et al.
 2024), LongVILA (Chen et al.
 2024), and Qwen2.5-VL (Bai et al.
-2025) as test models.
-To minimize any adverse effects on accuracy, all evaluations were performed using bf16 precision.
-As shown in Table 1, we compared the accuracy of Full-Attention, Block-Attention, APE, Star-Attention, and PEVLM in the datasets.
+2025) as representative models.
+To minimize accuracy loss, all experiments are conducted with bf16 precision.
+We compared the accuracy of Full-Attention, Block-Attention, APE, Star-Attention, and PEVLM in the datasets.
 Since the other methods are finetune-free, we did not finetune the weights in the Block Attention test.
-For APE, to accommodate the needs of large-scale deployment, we tested with T=0.
-Regarding the sink block (shared prefix), we selected the system prompt by referring to the code provided by APE.
-For Star-Attention, we implemented its equivalent algorithm on a single node, with the anchor size set equal to the block size, as the configuration in the Block-Attention paper (Ma, Wang, and Tian 2025).
-The block size was uniformly set to 4096 across all algorithms.
-Smaller block sizes resulted in reduced accuracy for all methods, while larger blocks significantly degraded inference performance.
-For PEVLM, we set the sink size and block size to about 16 frames, which corresponds to approximately 4k tokens for Qwen2.5-VL and LongVILA, and about 3k tokens for LLaVA-Video.
+For APE, we conducted experiments with a temperature setting of $T=1.0$.
+This choice was motivated by two considerations.
+First, lowering the temperature further leads to additional accuracy degradation.
+As previously observed, parallel encoding tends to produce lower attention weight distributions over context blocks in VLMs, rather than higher.
+Therefore, decreasing the temperature further skews the attention distribution away from that of Full-Attention, exacerbating the accuracy loss.
+Second, using $T=1.0$ is favorable for large-scale deployment.
+Regarding the sink block (shared prefix) size, we selected the system prompt by referring to the code provided by APE.
+For Star-Attention, we implemented its equivalent algorithm on a single node, with the anchor size set equal to the context block size, following the setup in the Block-Attention paper (Ma, Wang, and Tian 2025).
+All methods use a context block size of 4096 tokens.
+Smaller context block sizes resulted in reduced accuracy for all methods, while larger blocks significantly degraded inference performance.
+For PEVLM, we set the sink and context block size to 16 frames, which corresponds to approximately 4k tokens for Qwen2.5-VL and LongVILA, and about 3k tokens for LLaVA-Video.
 Results
-The experiments show that PEVLM achieves an accuracy comparable to, and in some cases even surpassing, that of Full-Attention across all three evaluated models.
-Both APE and Star-Attention also approach or exceed the accuracy of Full-Attention when Full-Attention’s position encoding is retained.
-However, Block-Attention exhibits a notable accuracy drop compared to Full-Attention in most settings, primarily due to the lack of alignment-specific fine-tuning.
-Although parallel encoding is generally expected to degrade performance, it can instead improve accuracy in long-context scenarios.
-This is because parallel encoding effectively shortens the total token length involved in each softmax computation, thereby mitigating the softmax feature degradation that occurs with long contexts (Veličković et al.
+As shown in Table 1, for shorter video scenarios, all methods achieve accuracy comparable to or even better than Full-Attention.
+Although parallel encoding is generally expected to degrade performance, it can actually improve accuracy in long-context scenarios.
+This is because the softmax function in attention mechanisms struggles to produce reliable outputs when the input length exceeds the training horizon (Veličković et al.
 2025).
-As a result, the softmax output more robustly approximates an ideal sharp function, which ultimately enhances the performance of some models under long-context conditions.
+However, parallel encoding effectively reduces the number of tokens involved in each softmax computation, thus alleviating the degradation of softmax behavior in long-context settings.
+As a result, the softmax output more robustly approximates an ideal sharp distribution, ultimately enhancing the performance of certain models under extended context lengths.
+While all parallel encoding methods perform well on short videos, their effectiveness declines as the temporal length of the video increases, especially when video durations extend to several or even tens of minutes.
+In such scenarios, PEVLM consistently outperforms other parallel encoding approaches, particularly on the Qwen2.5-VL model.
+Qwen2.5-VL uses 3D-MROPE and is thus more sensitive to temporal order in videos.
+As APE and Star-Attention reuse position IDs across different context blocks, which severely distort temporal features, resulting in meaningless or garbled output.
+In contrast, Block-Attention reencodes positional information, which improves its accuracy on Qwen2.5-VL.
+PEVLM further improves accuracy by more rigorously preserving sequential positional encodings and incorporating a shared sink block, leading to better results.
+For LongVILA and LLaVA-Video, which use standard RoPE, the aforementioned issues are less pronounced.
+Nonetheless, PEVLM still outperforms other parallel encoding methods in long-video scenarios.
 Performance Evaluation
 In this section, we evaluate the performance of PEVLM.
-To facilitate experimentation, we built PEVLM basied on the SGLang, which is a fast serving framework for large language models and vision-language models, widely used in cloud production deployment.
-Figure 4: We evaluate PEVLM under different block_size settings.
-The sink_size is set to (system prompt+one frame).
+Figure 4: PEVLM under different context_block_size settings.
 "Attn Time" refers to attention computation time, and "Other Time" covers all other LLM costs.
 "Attn SpeedUP" and "LLM SpeedUP" indicate the acceleration over Attention layer and overall LLM performance.
+The sink_size is set to (system prompt+one frame).
+Figure 5: PEVLM under different sink_size settings.
+"Attn Time" refers to attention computation time, and "Other Time" covers all other LLM costs.
+"Attn SpeedUP" and "LLM SpeedUP" indicate the acceleration over Attention layer and overall LLM performance.
+The context_block_size is set to (16 frames).
 Setup
-Since the LLM component in VLM models remains largely consistent across different implementations, performance differences arise primarily from hyperparameter configurations rather than model weights.
-Consequently, we conducted computational efficiency evaluations on the Qwen2.5-VL model.
+To facilitate experimentation, we built PEVLM based on the SGLang, which is a fast serving framework for large language models and vision-language models, widely used in cloud production deployment.
+We conducted computational efficiency evaluations on the Qwen2.5-VL model.
 The primary objective of this experiment is to assess the improvements in computational efficiency introduced by PEVLM.
@@
 Given that PEVLM primarily optimizes the attention mechanism within the LLM, we measured the execution time of both the entire LLM module and its individual attention layers separately.
-Figure 5: We evaluate PEVLM under different sink_size settings.
-The block_size is set to (16 frames).
-"Attn Time" refers to attention computation time, and "Other Time" covers all other LLM costs.
-"Attn SpeedUP" and "LLM SpeedUP" indicate the acceleration over Attention layer and overall LLM performance.
 Results
-As illustrated in Figure 4 and Figure 5, the end-to-end inference speed of the LLM is improved by 2.58× to 3.14×.
-Notably, the attention module achieves a speedup ranging from 7.79× to 33.81×, while the execution time of non-attention components remains largely unaffected.
-As the block size or sink size increases, the acceleration effect of the attention module decreases significantly.
-This observation is consistent with the Equation 11.
-Moreover, because non-attention components constitute a substantial portion of the total runtime in PEVLM, the overall end-to-end speedup of the LLM moderately decreases from 3.14× to 2.58×.
+As shown in Figure 4 and Figure 5, the end-to-end inference speed of the LLM improves by 2.58× to 3.14× as the block size and sink size vary.
+The attention module, in particular, achieves substantial acceleration, with speedups ranging from 7.79× to 33.81×, while the runtime of non-attention components remains largely stable.
+As the size of the sink block or context blocks increases, the speedup of the attention module correspondingly decreases.
+This trend is consistent with Equation 10, which indicates that the computational cost of attention grows quadratically with the size of the sink and context blocks.
+However, since non-attention components account for a considerable portion of total runtime, the overall end-to-end speedup degrades only moderately.
+However, as the context blocks size and sink block size increase to 16 frames, the proportion of attention in the total runtime increases from 6.4% to 22.9%, indicating that further increases in block size will lead to a more pronounced drop in overall performance.
+We will analyze the trade-off between performance degradation and potential accuracy gains from larger sink/context blocks size in the following sections.
 Accuracy with Limited Latency
-One key challenge in deploying VLMs on edge devices is the limited computational resources and strict latency requirements.
-To evaluate the acceleration performance of PEVLM under this scenario, we designed a simulation accordingly.
-A simplified experimental setup was implemented to assess the efficiency of PEVLM in edge deployment settings.
+Deploying VLMs on edge devices presents significant challenges due to constrained computational resources and stringent latency requirements.
+To evaluate the practical benefits of PEVLM in such settings, we designed a latency-aware simulation that mimics real-world deployment conditions on resource-limited devices.
+Setup
+We adopt LongVideoBench as the evaluation benchmark and introduce a latency threshold during inference: any sample that exceeds this latency limit is treated as a failure.
+All experiments are conducted on the NVIDIA H20-96G GPGPU.
+We use Qwen2.5-VL-7B-Instruct with Full-Attention as the primary baseline, and compare the performance of Full-Attention and PEVLM under identical latency constraints.
+To further reflect deployment in low-resource environments, where smaller models are commonly used to meet latency budgets, we also include comparisons with Qwen2.5-VL-3B-Instruct, a smaller variant Qwen2.5-VL-7B-Instruct.
 Table 2: Accuracy with Limited Latency
@@
 sink - - sys+2f sys+16f
-block - - 16f 16f
+context - - 16f 16f
 No Limit 60.43% 55.05% 61.18% 62.23%
@@
 20s 23.26% 24.38% 61.03% 60.28%
-Setup
-We use LongVideoBench as the evaluation benchmark and introduce a latency constraint during inference: any sample whose inference time exceeds the predefined latency threshold is considered a failure.
-The experiments are conducted on the H20-96B platform.
-We select Qwen2.5-VL-7B-Instruct with Full-Attention as the baseline, and compare Full-Attention, PEVLM, and using smaller models (Qwen2.5-VL-3B-Instruct).
 Results
-As shown in Table 2, the accuracy of Full-Attention drops significantly when the latency constraint is reduced from 40s to 30s, while the accuracy of Full-Attention 3B also declines sharply when the latency constraint is tightened from 30s to 20s.
-PEVLM maintain relatively high accuracy even at 20s.
-Notably, although larger sink and block sizes (e.g., sys+16frames, block=16frames) achieve higher accuracy under looser latency constraints, the smaller sink size (sink=sys+2frames, block=16frames) delivers better accuracy at the 20s latency limit due to its lower latency.
+As shown in Table 2, the accuracy of Full-Attention (7B) drops sharply when the latency constraint is reduced from 40s to 30s, and the smaller Full-Attention (3B) model also shows a significant decline when the limit is further tightened to 20s.
+In contrast, PEVLM maintains consistently high accuracy even under the strictest 20s constraint.
+Notably, while the larger sink size (sys+16f) achieves the best accuracy under relaxed constraints ($\geq$30s), the smaller configuration (sys+2f) performs better under the 20s limit due to its lower latency overhead.
+In general, PEVLM significantly improves accuracy in latency-constrained scenarios.
+Moreover, to maximize performance across diverse hardware and deployment conditions, the optimal configuration (e.g., sink/block size) should be adaptively tuned.
 How does each component in PEVLM contribute to the performance?
-In Table 3, we present an ablation study to evaluate the contribution of each component in PEVLM, including the use of Sequential Position Embedding (SP), Dividing the Video by Frames (DF), and Adding Frames to the Sink (FS).
-The results are averaged across the four base models reported in Table 1.
-We observe that sequential parallel encoding yields the most significant improvement in accuracy, achieving an average performance gain of 12.64%.
-Incorporating the initial frames into the sink block leads to an additional average accuracy gain of 1.54%.
-In contrast, dividing the video by frames provides only a marginal improvement compared to directly segmenting the video without this strategy.
 Table 3: Ablation study of PEVLM components on LongVideoBench
@@
 $\checkmark$ $\checkmark$ $\checkmark$ 58.26% 62.22%
-Position Encoding
-We further analyze the component with the greatest impact on accuracy: the use of sequential position embeddings.
-As discussed before, adopting sequential position embeddings instead of reusing position embeddings across blocks produces attention distributions that more closely resemble those of Full-Attention, thereby aligning better with the model’s expected behavior.
+In Table 3, we present an ablation study to evaluate the contribution of each component in PEVLM, including the use of Sequential Position Embedding (SP), Dividing the Video by Frames (DF), and Adding Frames to the Sink (FS).
+The results are averaged across the four base models reported in Table 1.
+We observe that sequential parallel encoding yields the most significant improvement in accuracy, achieving an average performance gain of 12.64%.
+Incorporating the initial frames into the sink block leads to an additional average accuracy gain of 1.54%.
+In contrast, dividing the video by frames provides only a marginal improvement compared to directly segmenting the video without this strategy.
+Sequential Position Embeddin
+We further analyze the component with the greatest impact on accuracy: the preservation of sequential position embeddings.
+As discussed before, adopting sequential position embeddings instead of reusing position embeddings across context blocks produces attention distributions that more closely resemble those of Full-Attention, thereby aligning better with the model’s expected behavior.
 As shown in Table 4, the use of sequential position embeddings significantly improves accuracy in both Qwen2.5-VL and LongVILA.
@@
 PEVLM 14.73% 62.23%
-APE (T=1.0) 51.46% 53.40%
+LongVILA- APE (T=1.0) 51.46% 53.40%
 7B-256f Star Attn 51.76% 53.63%
@@
 7B-Qwen2 Star Attn 58.71% 58.49%
-(128) PEVLM 58.34% 59.01%
-Sink & Block Size
-As analyzed in the third section, the inference latency of PEVLM increases quadratically with the sink size, which is consistent with the experimental results shown in Figure 5.
-Furthermore, we investigate the impact of sink size on model accuracy.
-As observed in the second section, it is essential to involve the initial frames of the video in the sink.
-This observation is validated in Figure 6, where incorporating the initial frames into the sink leads to varying degrees of accuracy improvement across different models.
+PEVLM 58.34% 59.01%
+Sink & Context Block Size
+As analyzed in the PEVLM section, the inference latency of PEVLM grows quadratically with the sizes of both the sink and context blocks, which aligns with the empirical results shown in Figure 4 and Figure 5.
+We further investigate how these sizes affect model accuracy.
+As noted in the Observations section, including the initial frames of the video in the sink is crucial.
+This is validated in Figure 6, where incorporating early frames into the sink leads to noticeable accuracy improvements across all evaluated models.
 Figure 6: The impact of sink sizes on accuracy
-Moreover, as shown in Figure 6, the accuracy of all models shows an upward trend as the sink size increases.
-However, because of the quadratic increase in latency, using longer sink sizes is not favorable for performance.
-Therefore, it may be necessary to conduct a simple test prior to deployment to determine an appropriate sink size.
-For the three models shown in Figure 6, setting the sink size between 14 and 18 already yields strong accuracy.
-If latency constraints are also taken into account, this could lead to a trade-off between performance and accuracy.
-For example, as demonstrated in the results of Table 2, under a 20-second latency constraint, using only sys+2frames as the Sink Block achieves better accuracy than using sys+16frames.
+Figure 6 also shows that model accuracy generally increases with sink size.
+However, due to the quadratic latency growth, excessively large sink sizes are not ideal for deployment.
+A simple pre-deployment test may be necessary to determine the optimal configuration.
+For the three models in Figure 6, a sink size between 14 and 18 already achieves strong performance.
+When latency constraints are considered, this leads to a trade-off between accuracy and efficiency.
+For example, as shown in Table 2, under a 20-second latency budget, using only [SYS] + 2 frames as the Sink Block yields higher accuracy than [SYS] + 16 frames.
 Figure 7: The impact of block sizes on accuracy
-As shown in Figure 4, the inference latency of PEVLM also increases quadratically with block size, which is consistent with the experimental results shown in Figure 4.
-However, the block size has a more significant impact on performance because the coefficient of the block size’s quadratic term in Equation 11 is much larger than that of the sink size.
-We investigate the impact of block size on model accuracy here.
-As shown in Figure 7, the model accuracy exhibits an upward trend as the block size increases, similar to sink size.
-However, due to the quadratic growth in latency, a trade-off arises between computational efficiency and accuracy.
+We also examine the effect of context block size on model accuracy.
+As shown in Figure 7, increasing the size of context blocks leads to consistent accuracy gains, similar to the sink size.
+However, Figure 4 reveals that latency also increases quadratically with block size, highlighting the same trade-off between accuracy and computational efficiency.
 Conclusion
-This study investigates the potential of parallel-encoding-based methods in VLMs, which can reduce the computational complexity of attention mechanisms for fast inference and reuse positions for long-context processing but often result in performance degradation.
-To address this issue, we propose PEVLM, a training-free method designed to enable accurate, fast, and long-context video processing.
-PEVLM achieves this by aligning the attention weight distribution of parallel encoding with Full-Attention through two key steps: selecting an appropriate sink size and block size, and utilizing sequential position embeddings.
-Empirically, we demonstrate that PEVLM enhances both accuracy and efficiency in long video processing tasks while successfully scaling to process hundreds of chunks in parallel across various settings.
+We introduce PEVLM, a fine-tuning-free parallel encoding strategy that significantly accelerates Vision-Language Models for long-video understanding.
+By preserving sequential position embeddings and leveraging a shared Sink Block, PEVLM aligns attention behavior with Full-Attention while reducing complexity from $O((T\times N)^{2})$ to $O(T\times N)$.
+It achieves up to 7.47$\times$ attention speedup, 40% end-to-end latency reduction, and strong accuracy gains under tight latency constraints.
+PEVLM offers a practical and scalable solution for real-world, long-context multimodal tasks.
 Future Work
@@
 Such an extension would enable real-time video-language reasoning under strict latency constraints and dynamic, continuously evolving environments.
-Appendix A Benchmark Details
+One additional direction for future exploration is to investigate why parallel encoding methods can surpass Full-Attention in accuracy.
+Our preliminary analysis suggests that this may be due to the shortened input length in the attention softmax, which improves attention sharpness and reliability in long-context settings.
+However, since this work focuses on applying and optimizing parallel encoding for VLMs, we leave a deeper theoretical and empirical investigation of this phenomenon for future work.
+Understanding its root cause may offer new insights into further boosting model performance on long contexts.
+Appendix A. Benchmark Details
 We evaluate our method on several video understanding benchmarks that test different aspects of video comprehension:
@@
 It aims to assess whether models can maintain coherence, memory, and attention over prolonged contexts, making it a rigorous testbed for long-form video modeling.
-Appendix B Attention Weights Distribution of LLaVA-Video and LongVILA
+Appendix B. Attention Weights Distribution of LLaVA-Video and LongVILA
 (a) Full-Attention